### v1

# Simultaneous Motion And Noise Estimation with Event Cameras

Abstract
Event cameras are emerging vision sensors, whose noise is challenging to characterize. Existing denoising methods for event cameras consider other tasks such as motion estimation separately (i.e., sequentially after denoising). However, motion is an intrinsic part of event data, since scene edges cannot be sensed without motion. This work proposes, to the best of our knowledge, the first method that simultaneously estimates motion in its various forms (e.g., ego-motion, optical flow) and noise. The method is flexible, as it allows replacing the 1-step motion estimation of the widely-used Contrast Maximization framework with any other motion estimator, such as deep neural networks. The experiments show that the proposed method achieves state-of-the-art results on the E-MLB denoising benchmark and competitive results on the DND21 benchmark, while showing its efficacy on motion estimation and intensity reconstruction tasks. We believe that the proposed approach contributes to strengthening the theory of event-data denoising, as well as impacting practical denoising use-cases, as we release the code upon acceptance. Project page: https://github.com/tub-rip/ESMD
1 Introduction
Event cameras are emerging vision sensors that overcome challenges of conventional cameras (e.g., motion blur, limited dynamic range, data and power efficiency), yet they suffer from a considerable amount of noise due to their novelty and operation in low-power (transistor subthreshold) conditions [19]. Since event cameras are suitable for many computer vision tasks, especially motion-related tasks, it is paramount to classify event data that is related to motion (i.e., signal) and that is not (i.e., noise). However, denoising event data is a challenging problem, since noise properties are not fully characterized to date, and it is not feasible to define ground-truth (GT) noise labels in real-world data recordings. Previous work either ($i$) leveraged simulation to generate pure signal (events from edge motion under constant illumination) and noise, or ($ii$) prepared real event data that was expected to be clean, by aggressively removing noise-like events with existing filters and using bright illumination that minimized noise occurrence.
Figure 1: Overview of the proposed method. We use only raw events as input, and estimate signal events, noise events, and motion (e.g., dense optical flow and ego-motion estimation).
However, such artificial modifications on the real event data could change event signal/noise characteristics, possibly deteriorating the information that the original event had for downstream tasks (e.g., motion estimation). Our work is inspired by the observation that motion is an intrinsic part of event data since scene edges cannot be sensed without motion. Rather than solving denoising and motion estimation tasks separately, our idea is to leverage motion information to improve denoising (as “vice-versa” is well known). In this work (Fig. 1), we rethink and investigate the connection between denoising and motion estimation, by asking ourselves whether they can be combined in a more integrated way to benefit the results of both tasks simultaneously. We formulate the simultaneous estimation problem by leveraging the idea of motion compensation in the Contrast Maximization (CMax) framework, defining novel metrics about the contribution to the contrast score for each event. In summary, we present several distinctive contributions:
  1. 1.
    We propose a novel method to jointly estimate noise and motion only from event data (Sec. 3).
  2. 2.
    We demonstrate that our joint estimation method ($i$) achieves state-of-the-art results in existing denoising benchmarks (Sec. 4.2), and ($ii$) improves the robustness of motion estimation (CMax framework) (Sec. 4.3).
  3. 3.
    We show the flexibility of our method on ($i$) further applications, such as intensity reconstruction, validating its image quality (Secs. 4.4 and 4.5), and ($ii$) in combination with learning-based motion estimators (Sec. 4.6).
We hope our work opens a new line of research for event cameras, by considering the problems of motion estimation and denoising together to take advantage of their coaction.
2 Related Work
Denoising is a fundamental problem in event cameras since there are a lot of leakage events [24], and event-processing algorithms may suffer in scenarios with high noise [3]. Removing event noise has a double effect: it reduces the load of subsequent processing stages (as the load is proportional to the number of events) and improves their output (as the input data becomes of higher quality). In this work, we focus on noise due to the leakage inside the camera hardware, such as background activity (BA) events [35]. Events due to flickering or active lights can be treated as “noise” for methods that consider passive sensing, which are a large majority in computer vision, and are therefore not our focus.
Classical denoising works focus on the spatio-temporal correlation of events, as BA events have fewer neighboring events in space-time coordinates than signal events. This approach includes widely-used BA Filter (BAF) [9], and other types of spatio-temporal filters [36, 30, 56, 26]. Some works utilize additional sensor inputs (e.g., frames) to estimate noise probabilities [4, 12, 55, 29, 57], since edges are informative of the signal events produced.
Recently, learning-based approaches have gained more attention, such as multilayer perceptrons [24], convolutional neural networks [4, 12, 1], graph neural networks [2], and transformers [28]. They provide the probability of noise (i.e., signal-noise classification). Nonetheless, the challenge of learning-based approaches is the need for ground truth (GT) noise labels for training. GT labels have been obtained through simulation (e.g., [13, 24]) as well as real-world recordings, however, real-world recordings aggressively remove noisy events using existing filters (e.g., [24]) or use unnatural illumination settings (e.g., [28]), by leveraging the differences of noise occurrence due to different scene brightness [24, 34]. While these approaches are effective for benchmarking, the challenge of obtaining GT labels for practical real-world datasets remains unsolved.
Our work builds upon a series of successful methods on motion estimation and their applications. Event-based motion estimation is a paramount topic since event cameras naturally respond to motion in the scenes [19]. In particular, Contrast Maximization (CMax) [17, 49] is a state-of-the-art framework for motion estimation, which has various applications, such as rotational motion [15, 23, 44, 41, 31], optical flow [62, 43, 47, 27], intensity reconstruction [58], and SLAM [25, 22, 54]. We investigate the connection and potential benefits of applying the CMax framework to denoising during the estimation process (i.e., simultaneously), demonstrating motion estimation improvements as well as competitive denoising accuracy. A prior work, “ProgressiveMotionSeg” [8], has extended the CMax-based segmentation method [50] with extra weights that accounted for how much each event contributed to estimated motion clusters. However, [8] mainly focuses on the task of motion segmentation using low-DOF warps (e.g., 2-DOF feature flow) and does not discuss denoising efficacy or evaluate on standard denoising benchmarks. In contrast, our method supports various types of motion estimation (both low-DOF and high-DOF warps like optical flow), and considers that noise shall not be used to estimate motion, as from first principles we deem noise and motion to be uncorrelated. We propose signal-noise classifiers and show ranking invariance, which results in state-of-the-art denoising accuracy and improvements on motion estimation from noisy data.
3 Methodology
The following sections present the CMax framework (Sec. 3.1), the proposed signal/noise classifier (Sec. 3.2), and the simultaneous denoiser and motion estimator (Sec. 3.3).
Event cameras consist of independent pixels that generate asynchronous “events” when the logarithmic brightness at the pixel increases or decreases by a predefined contrast sensitivity. Each event $e_{k}\doteq(\mathbf{x}_{k},t_{k},p_{k})$ contains the pixel coordinates $\mathbf{x}_{k}$, the timestamp $t_{k}$, and polarity $p_{k}\in\{+1,-1\}$ of the brightness change. Events occur asynchronously and sparsely on the pixel grid, resulting in a variable data rate based on the scene texture and dynamics.
3.1 Contrast Maximization
The CMax framework [17] assumes events $\mathcal{E}\doteq\{e_{k}\}_{k=1}^{N_{e}}$ are caused by moving edges, and it transforms their coordinates according to a motion model $\mathbf{W}$, producing a set of warped events $\mathcal{E}^{\prime}_{t_{\text{ref}}}\doteq\{e^{\prime}_{k}\}_{k=1}^{N_{e}}$ at a reference time $t_{\text{ref}}$:
$\vspace{-0.5ex}e_{k}\doteq(\mathbf{x}_{k},t_{k},p_{k})\;\,\mapsto\;\,e^{\prime}_{k}\doteq(\mathbf{x}^{\prime}_{k},t_{\text{ref}},p_{k}).\vspace{-0.5ex}$ (1)
The warp $\mathbf{x}^{\prime}_{k}=\mathbf{W}(\mathbf{x}_{k},t_{k};\boldsymbol{\theta})$ transports each event from $t_{k}$ to $t_{\text{ref}}$ along the motion curve that passes through it. Then, they are aggregated on an image of warped events (IWE):
$\vspace{-0.5ex}\textstyle I(\mathbf{x};\mathcal{E}^{\prime}_{t_{\text{ref}}},\boldsymbol{\theta})\doteq\sum_{k=1}^{N_{e}}\delta(\mathbf{x}-\mathbf{x}^{\prime}_{k}),\vspace{-0.5ex}$ (2)
where each pixel $\mathbf{x}$ sums the number of warped events $\mathbf{x}^{\prime}_{k}$ that fall within it. The Dirac delta is approximated by a Gaussian, $\delta(\mathbf{x}-\boldsymbol{\mu})\approx\mathcal{N}(\mathbf{x};\boldsymbol{\mu},\epsilon^{2}\mathtt{Id})$ with $\epsilon=1$ pixel.
GasTank (E-MLB) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
LEGO (E-MLB)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Beam (E-MLB)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Driving (DND21)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(a) Input (Raw Events) (b) IWE (No Warp) (c) Output (Signal Events) (d) IWE (Signal Events) (e) Estimated Flow (f) Reference Frame
Figure 2: Qualitative result of denoising and flow estimation on E-MLB and DND21 data. The driving sequence has $5$-Hz artificial noise. (b) is an alternative visualization of (a) with the same grayscale coding as (d) for ease of comparison. Flow color is given in Fig. 1.
[…]
Finally, an optimization algorithm iterates the above steps to find the motion parameters $\boldsymbol{\theta}^{\ast}$ that maximize the alignment of events caused by the same scene edge. Event alignment is measured by the strength of the edges of the IWE, which is related to image contrast [18].
Motion models. As exemplary warps $\mathbf{W}$ we focus on two popular problem settings: rotational motion (3-DOF angular velocity estimation [15, 44, 41, 23]) and dense optical flow (pixel-wise velocity estimation [62, 42, 47, 49], resulting in $2N_{p}$-DOFs, where $N_{p}$ is the number of pixels). For the angular velocity estimation during a small time interval, the warp is parametrized by the angular velocity $\boldsymbol{\theta}\equiv\boldsymbol{\omega}=(\omega_{x},\omega_{y},\omega_{z})^{\top}$, and $\mathbf{x}^{h}(t)\sim\mathtt{R}(\boldsymbol{\omega}t)\,\mathbf{x}^{h}(0)$, using calibrated homogeneous coordinates $\mathbf{x}^{h}$ and exponential coordinates $\mathtt{R}(\boldsymbol{\phi})\doteq\exp(\boldsymbol{\phi}^{\wedge})$ [6, 16]. The warp for dense optical flow is $\mathbf{x}^{\prime}_{k}=\mathbf{x}_{k}+(t_{k}-t_{\text{ref}})\,\mathbf{v}(\mathbf{x}_{k})$, where $\boldsymbol{\theta}=\{\mathbf{v}(\mathbf{x})\}_{\mathbf{x}\in\Omega}$ is a flow field on the image plane $\Omega$.
3.2 Denoising Based on Local Contrast
As Fig. 1 shows, our method is iterative. In each iteration, motion estimation is refined and events are classified into signal $\mathcal{E}_{s}$ or noise $\mathcal{E}_{n}$. Without labels or additional information, events rely on each other to realize this classification. To this end, we compute a score $c_{k}$ for each event, rank the events according to their score, and threshold at some value $T$. The threshold is computed by assuming the amount or ratio of noisy events is known $\eta$, i.e., $T\equiv T(\eta)$, which allows us to have high-level control over the denoising process. Hence we classify events into signal (the $\tau\doteq(1-\eta)$ percentage of events with highest scores) and noise:
$\begin{split}\mathcal{E}_{s}\equiv\mathcal{E}_{\text{signal}}&\doteq\{e_{k}\in\mathcal{E}\;|\;c_{k}>T(\eta)\},\\ \mathcal{E}_{n}\equiv\mathcal{E}_{\text{noise}}&\doteq\mathcal{E}\backslash\mathcal{E}_{\text{signal}}\quad(\text{complement of }\mathcal{E}_{\text{signal}}).\end{split}$ (3)
We studied different choices for the per-event score $c_{k}$ based on the agreement among warped events $\mathcal{E}^{\prime}=\mathcal{E}^{\prime}_{s}\cup\mathcal{E}^{\prime}_{n}$, through their IWEs, and concluded on two. Letting $I^{s}_{k}\equiv I^{s}_{k}(\mathbf{x}^{\prime}_{k}),I^{n}_{k}\equiv I^{n}_{k}(\mathbf{x}^{\prime}_{k})$ and $I_{k}\equiv I(\mathbf{x}^{\prime}_{k})=I^{s}_{k}+I^{n}_{k}$ be the IWE values of previously classified signal events, noise events and all events, respectively, the two choices are: ($i$) ${}^{1}c_{k}\doteq I_{k}$, and ($ii$) the ratio of signal events ${}^{2}c_{k}\doteq I^{s}_{k}/I_{k}$ (probabilistic viewpoint). In fact, both are equivalent: rule ${}^{1}c_{k}>T_{1}$ is the same as rule ${}^{2}c_{k}>T_{2}$ if $T_{2}=1-I^{n}_{k}/T_{1}$ and the event noise $I^{n}_{k}$ is known. We use ${}^{1}c_{k}$, the number of warped events (also called local contrast [18, 50]), because it is the simplest to interpret.
The intuition is that “signal” events (e.g., due to motion) warped according to the true motion lie all at motion-compensated edges, and therefore produce a sharp IWE. Noise events, produced all over the image plane, are not expected to contribute to the IWE’s sharpness. Hence, we may classify events into signal or noise according to how much they contribute to edge strength. The IWE (2) provides one such measure of edge strength: the higher $I(\mathbf{x})$, the more events are warped to that pixel location $\mathbf{x}$ (i.e., more events support the same scene edge), thus producing a sharper IWE. Hence, we define how much an event $e_{k}$ contributes to edge strength by means of the IWE (2), ${}^{1}c_{k}$.
Invariance. Note that (3) is invariant to monotonically increasing functions of $c_{k}$, as such functions preserve the order. The same $\mathcal{E}_{\text{signal}}$ is obtained if one ranks and thresholds (at the corresponding level) values $\{c_{k}\}$, $\{\ln(c_{k})\}$, $\{e^{c_{k}}\}$, etc. They are equivalent under the ranking rule (3).
Robustness to various edge strengths. The Gaussian kernel in the IWE controls the sensitivity to edge strength. Since the proposed method is based on both signal and noise IWEs, increasing the size of the Gaussian kernel in the IWE can emphasize contour edges over isolated points. This preserves signal events better in regions with low IWE intensities, for example at pixels with large depth values.
3.3 Alternating Optimization
Classifying events into signal and noise requires knowledge of the true motion, and estimation of the true motion (e.g., using CMax) requires knowledge of the signal events because, by definition, noise events carry no information about motion. Hence, it is a chicken-and-egg problem.
We approach its solution iteratively, as shown in Fig. 1, by combining motion estimation (Sec. 3.1) and S/N classification (Sec. 3.2). Assuming some initialization, the current signal events $\mathcal{E}_{\text{signal}}$ are used to estimate the motion via CMax (1 step is enough); then all events are warped to compute the IWE and the per-event scores $c_{k}$, from which the sets (3) are recomputed, in preparation for the next iteration.
In practice, we initialize with a random split of $\mathcal{E}$ into $\mathcal{E}_{\text{signal}}$ and $\mathcal{E}_{\text{noise}}$, update $\mathcal{E}_{\text{signal}}$ at every warp during the optimization iteration until the process converges (as flagged by the convergence of the motion parameters in CMax).
Computational complexity. The computational complexity of one iteration of the proposed method is $O(N_{p}+N_{e}\log N_{e})$, which is slightly larger than the original CMax (i.e., $O(N_{p}+N_{e})$) due to the search for the highest-ranked events during the optimization.
4 Experiments
We evaluate the proposed method on four datasets using various metrics (in Sec. 4.1). In the following sections using common benchmarks, we assess denoising (Sec. 4.2), joint motion estimation (Sec. 4.3), and the efficacy on intensity reconstruction (Sec. 4.4). We also conduct sensitivity analyses (Sec. 4.5) and an ablation (Sec. 4.6).
E-MLB (Day) E-MLB (Night) DND21 ND1 ND4 ND16 ND64 ND1 ND4 ND16 ND64 Model-based Raw $0.821$ $0.824$ $0.815$ $0.786$ $0.89$ $0.824$ $0.786$ $0.768$ $0.869$ BAF [9] $0.861$ $0.869$ $0.876$ $0.89$ $0.946$ $0.973$ $0.992$ $0.942$ $0.92$ KNoise [30] $0.846$ $0.837$ $0.83$ $0.807$ $0.954$ $0.956$ $0.871$ $0.817$ $0.887$ DWF [24] $0.878$ $0.876$ $0.866$ $0.865$ $0.923$ $0.962$ $0.988$ $0.932$ $0.905$ EvFlow [53] $0.848$ $0.878$ $0.868$ $0.833$ $0.969$ $0.983$ $0.889$ $0.797$ $1.006$ Ynoise [14] $0.866$ $0.863$ $0.857$ $0.821$ $1.009$ $0.943$ $0.875$ $0.792$ $0.966$ TS [33] $0.877$ $0.887$ $0.87$ $0.837$ $1.033$ $0.944$ $0.886$ $0.797$ $0.985$ IETS [5] $0.772$ $0.785$ $0.777$ $0.753$ $0.950$ $0.823$ $0.804$ $0.711$ $0.900$ GEF [11] $1.051$ $0.938$ $0.935$ $0.927$ $1.027$ $0.955$ $0.946$ $0.935$ $0.932$ Ours 0.938 254 384 292 974 70.93825438429297470.938\,254\,384\,292\,974\,70.938 254 384 292 974 7 0.958 464 335 277 024 80.95846433527702480.958\,464\,335\,277\,024\,80.958 464 335 277 024 8 0.985 896 450 122 991 80.98589645012299180.985\,896\,450\,122\,991\,80.985 896 450 122 991 8 0.950 081 880 240 459 60.95008188024045960.950\,081\,880\,240\,459\,60.950 081 880 240 459 6 1.036 831 015 641 0911.0368310156410911.036\,831\,015\,641\,0911.036 831 015 641 091 0.961 426 679 598 1070.9614266795981070.961\,426\,679\,598\,1070.961 426 679 598 107 0.944 527 953 022 538 40.94452795302253840.944\,527\,953\,022\,538\,40.944 527 953 022 538 4 0.932 362 037 844 1260.9323620378441260.932\,362\,037\,844\,1260.932 362 037 844 126 0.992 069 121 10.99206912110.992\,069\,121\,10.992 069 121 1 Learning EventZoom [12] $0.996$ $0.988$ $0.996$ $0.97$ $1.055$ $1.007$ $1.01$ $0.988$ $1.059$ EDnCNN [4] $0.887$ $0.908$ $0.903$ $0.912$ $1.001$ $1.024$ $1.079$ $1.086$ $0.977$ MLPF [24] $0.851$ $0.855$ $0.846$ $0.84$ $0.926$ $0.928$ $0.91$ $0.906$ $0.944$ EDformer [28] $0.952$ $0.955$ $0.956$ $0.942$ $1.048$ $1.019$ $1.076$ $1.099$ $1.041$
[…]
1Hz 5Hz 10Hz hotel driving hotel driving hotel driving Model-based KNoise [30] $0.6773$ $0.6296$ $0.6703$ $0.6235$ $0.6413$ $0.6142$ DWF [24] $0.9268$ $0.7409$ $0.8620$ $0.6901$ $0.7958$ $0.6563$ BAF [9] $0.9535$ $0.8479$ $0.8916$ $0.7930$ $0.8366$ $0.7479$ Ynoise [14] $0.9690$ $0.9409$ $0.9234$ $0.9093$ $0.8987$ $0.8800$ TS [33] $0.9716$ $0.9307$ $0.9606$ $0.9270$ $0.9620$ $0.9202$ Ours 1.013 802 194 317 176 21.01380219431717621.013\,802\,194\,317\,176\,21.013 802 194 317 176 2 0.882 119 693 133 694 80.88211969313369480.882\,119\,693\,133\,694\,80.882 119 693 133 694 8 0.962 515 815 619 444 40.96251581561944440.962\,515\,815\,619\,444\,40.962 515 815 619 444 4 0.855 398 484 638 116 20.85539848463811620.855\,398\,484\,638\,116\,20.855 398 484 638 116 2 0.960 564 724 999 644 20.96056472499964420.960\,564\,724\,999\,644\,20.960 564 724 999 644 2 0.836 273 555 102 2360.8362735551022360.836\,273\,555\,102\,2360.836 273 555 102 236 Learning EDnCNN [4] $0.9573$ $0.8873$ $0.9365$ $0.8748$ $0.9006$ $0.874$ MLPF [24] $0.9704$ $0.8887$ $0.9704$ $0.8845$ $0.9634$ $0.8761$ EDformer [28] $0.9928$ $0.9541$ $0.9845$ $0.9424$ $0.9699$ $0.9264$
[…]
Figure 3: ROC curves on the DND 21 dataset sequences. Dash lines represent learning-based methods (see also Tab. 2).
Figure 4: Comparison of various denoising methods on the two DND21 sequences (driving and hotel) at two different noise levels.
4.1 Datasets, Metrics, and Hyper-parameters
Datasets. E-MLB [10] is a large-scale, de-facto dataset to benchmark denoising. It consists of $100$ sequences with 4 different brightness levels using neutral-density (ND) filters on a DAVIS346 camera ($346\times 260$ px) during day and night.
Historically, the two sequences (hotel and driving) of the DND21 dataset [24] have been widely used for denoising evaluation [24, 10, 28]. Both sequences are recorded with a DAVIS346 camera [52]. They are regarded as “signal” data, as the sequences have been aggressively filtered offline. Combined with other pure BA noise sequences, the dataset has event-wise noise annotations that are useful for evaluation. We use the above sequences with different noise rates, $1,3,5,7,10$ Hz per pixel, following prior work [28].
The ECD dataset [39] is a standard dataset for various tasks including camera ego-motion estimation [15, 60, 46, 40, 59, 48]. Using a DAVIS240C camera ($240\times 180$ px [7]), each sequence provides events, frames, calibration information, IMU data, and ground truth (GT) camera poses (at $200$ Hz). We use boxes_rotation and dynamic_rotation sequences to evaluate rotational motion estimation accuracy.
Finally, we also use sequences from MVSEC [61] and zurich_city_12a sequence from the DSEC dataset [20], which are typically used for optical flow estimation. We use MVSEC sequences for quantitative evaluation and the DSEC sequence for qualitative evaluation, since it is captured during the night and hence includes lots of noise.
Metrics. The metrics used for denoising assessment are: receiver operating characteristic (ROC) curve, area under the curve (AUC) and Mean Event Structural Ratio (MESR). Given signal and noise GT labels, the ROC shows the True-Positive Rate and False-Positive Rate over several S/N ratios. The AUC of the ROC becomes closer to $1$, the better the denoiser works. The MESR is a denoising metric that does not require GT labels [10]. We follow previous work to fix the number of signal events to 30000 [28, 10], as well as $M=20000$ [10, Eq.10], to discount the difference in the number of signal events from various denoising methods. The higher, the better.
We use the RMS error for angular velocity estimation and the EPE for optical flow estimation, following previous works [41, 23, 25, 49], as well as the FWL metric [51], which is the relative variance (sharpness) of the IWE with respect to that of the identity warp.
Hyper-parameters. For the denoising experiments, we test various values $\tau=\{0.9,\ldots,0.1\}$ to calculate the ROC on DND21 and $\tau=\{0.9,\ldots,0.7\}$ for the RMS on ECD. For E-MLB benchmarking, we fix the number of signal events to follow the prior work [10]. To analyze the proposed pipeline in the CMax framework, we use model-based rotational motion estimation [15] and tile-based optical flow estimation [47] approaches. We use the magnitude of the IWE gradient [18] as the CMax objective function.
4.2 Denoising Results
We evaluate denoising accuracy using AUC on DND21 data, and using MESR on E-MLB and DND21 data, using the optical flow estimation.
Denoising Performance on E-MLB. Table 1 shows the benchmark results on the E-MLB dataset using the MESR metric. We categorize prior work into model-based methods and learning-based methods for convenience. Compared with other model-based techniques, our method ranks first or second, showcasing the efficacy of the proposed method on denoising. Even compared with learning-based ones, our method achieves competitive scores (e.g., second in E-MLB (Day) ND4–ND64 conditions, etc). Notice that, however, learning-based approaches utilize additional information (GT labels) during training, while our method does not. Nonetheless, the results demonstrate the efficacy of our approach, which contributes to the theoretical foundation for the denoising methodology of event data. We discuss further details of the evaluation metrics (ESR) in Sec. 7.
Qualitative results are shown in Fig. 2. The denoised outputs (“Signal”) are reasonable, as dark (underexposed) regions of the scene become cleaner, and the IWEs using the estimated flow provide sharp edges of the original scenes. We compare with state-of-the-art methods in detail next.
Refer to caption Refer to caption Refer to caption
Raw Events IWE(0) GT rotation
IWE of signal
Refer to caption Refer to caption Refer to caption
Estimated rotation
Refer to caption Refer to caption Refer to caption
Signal events
Refer to caption Refer to caption Refer to caption
Noise events
Refer to caption Refer to caption Refer to caption
(a) Random sampling (b) BA filter (BAF) [9] (c) Ours
[…]
Denoising Accuracy on DND21. Table 2 shows the result of AUC for the ROC result on the DND21 dataset. Our method achieves consistently high AUC compared with existing model-based methods. Note that the DND21 dataset aggressively removes events to generate “pure-signal” sequences [24, Sec.5]. Due to this step, the labeled signal events may be fewer than the actual ones corresponding to the true motion. Hence, the DND21 sequences could result in a slight degradation in the False Positive score for motion-based (unsupervised) denoising methods like ours, which may result in lower AUC values. Nonetheless, our method provides competitive AUC scores among other state-of-the-art. The ROC curves are shown in Fig. 3.
Figure 6: Results on the ego-motion estimation application (top: boxes, bottom: dynamic) [39]. Our method relaxes the dependency on the initial value for CMax-based rotational motion estimation (indicated as “Gain”).
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(a) Input events (b) Estim. flow (c) Signal events (d) Reconstr. from (c) signal (e) Reconstr. from (a) input (f) Reference frame
Figure 7: Image intensity reconstruction on E-MLB data [10]. The first two rows use the reconstruction method EVILIP [58] and the last two rows use E2VID [45]. Reconstruction on denoised events produces higher quality images (i.e., fewer artifacts) than on raw event data.
dynamic_rot boxes_rot RMS $\downarrow$ FWL $\uparrow$ RMS $\downarrow$ FWL $\uparrow$ $5$ Hz Noise CMax [17] 20.000 709 6820.0007096820.000\,709\,6820.000 709 68 1.258 796 0971.2587960971.258\,796\,0971.258 796 097 124.641 377124.641377124.641\,377124.641 377 1.128 529 5931.1285295931.128\,529\,5931.128 529 593 – w/ Init. 8.274 876 2458.2748762458.274\,876\,2458.274 876 245 1.273 697 0451.2736970451.273\,697\,0451.273 697 045 20.659 240 3320.6592403320.659\,240\,3320.659 240 33 1.214 159 4591.2141594591.214\,159\,4591.214 159 459 Downsampling (Best) 8.807 703 958.807703958.807\,703\,958.807 703 95 1.272 907 8221.2729078221.272\,907\,8221.272 907 822 124.868 639 7124.8686397124.868\,639\,7124.868 639 7 1.107 758 3951.1077583951.107\,758\,3951.107 758 395 – w/ Init. 8.226 219 4448.2262194448.226\,219\,4448.226 219 444 1.273 725 3071.2737253071.273\,725\,3071.273 725 307 20.618 615 8620.6186158620.618\,615\,8620.618 615 86 1.214 188 4551.2141884551.214\,188\,4551.214 188 455 Ours (Best) 8.521 574 0738.5215740738.521\,574\,0738.521 574 073 1.273 446 2961.2734462961.273\,446\,2961.273 446 296 97.584 574 8497.5845748497.584\,574\,8497.584 574 84 1.140 880 661.140880661.140\,880\,661.140 880 66 – w/ Init. $8.170$ 1.273 725 2991.2737252991.273\,725\,2991.273 725 299 20.604 334 6220.6043346220.604\,334\,6220.604 334 62 1.214 179 0181.2141790181.214\,179\,0181.214 179 018 BAF 19.674 893 0519.6748930519.674\,893\,0519.674 893 05 1.259 772 491.259772491.259\,772\,491.259 772 49 125.028 145 1125.0281451125.028\,145\,1125.028 145 1 1.127 308 9021.1273089021.127\,308\,9021.127 308 902 – w/ Init. 8.253 078 2028.2530782028.253\,078\,2028.253 078 202 1.273 636 991.273636991.273\,636\,991.273 636 99 20.550 436 620.550436620.550\,436\,620.550 436 6 1.213 505 481.213505481.213\,505\,481.213 505 48
Table 3: Angular velocity estimation on ECD dataset [39].
Qualitative comparisons among several denoising methods is given in Fig. 4. From the results, our method achieves a good trade-off between under-denoising (preserving not only scene-edge details but some noise) and over-denoising (removing not only most of the noise but some edge details). This can be confirmed, for example, in the upper-right edges in the driving example, and the person’s arm in the hotel example. Essentially, all denoising methods have some parameters to control the amount of denoising; our proposed method has the signal ratio, which is intuitive; other optimization methods have some parameters about spatio-temporal kernel size, and learning-based approaches use the probability threshold for the network output. While we fix it for each dataset and condition for fair comparison (we do not tune it for each sequence in the results), this is an essential challenge for denoising algorithms in general, regardless of their underlying character (learning-based or model-based). Hence, we believe it is important for research to discuss denoising effectiveness in combination with downstream tasks, as shown in Secs. 4.3 and 4.4.
4.3 Joint Motion Estimation Results
Ego-motion on ECD. One of the merits of the proposed joint estimation method is that it can improve motion estimation results. Here, the results of ego-motion (e.g., angular velocity) estimation are displayed in Fig. 5. We used the classical ECD dataset and injected BA noise from the DND21 recordings with different noise rates. As no prior work simultaneously estimates motion and noise, we compare against sequential filters: random sampling (of the input data with a uniform distribution) and BAF [9]. As shown in the IWEs, our method produces the sharpest edges (cf. (b) and (c)) and the most reasonable signal-noise classification among others (cf. (a), (b), and (c)). Note that the original ECD recordings have noise events besides the ones that are injected using DND21 recordings, which makes the calculation of ROC (similar to Sec. 3.2) intricate. Nonetheless, we report ROC results in the supplementary.
The quantitative evaluation on angular velocity estimation is reported in Fig. 6, as well as the summary of the supplementary detailed table in Tab. 3. As shown in the first row (“CMax”), the original CMax degrades due to the noise occurrences. Our method achieves the best results in terms of RMS and FWL, compared with the random sampling method as well as the original CMax (no denoising). Also, using better initialization (the result of the previous slice for sequential estimation) always improves the results (“w/ Init.”): the estimation accuracy relies on both the initialization and the denoising strategies. Hence, the result demonstrates that the proposed method relaxes the dependency on the initialization, thus improving the robustness of motion estimation.
Optical flow on MVSEC. The results on more complex motion estimation (e.g., optical flow) are shown in Tab. 4, using the MVSEC dataset [61], with BA noise from the DND21 dataset. Due to the $5$Hz BA noise, the baseline method MultiCM [47] degrades. Our denoising method improves the flow accuracy, showcasing the efficacy of simultaneous noise estimation. A challenge of evaluating denoising in these real-world datasets (e.g., ECD, MVSEC) is the existing noise in the original sequences. Hence, we only evaluate motion estimation metrics, not ROC or AUC.
indoor1 indoor2 EPE $\downarrow$ %Out $\downarrow$ EPE $\downarrow$ %Out $\downarrow$ MultiCM [47] 2.675 668 1172.6756681172.675\,668\,1172.675 668 117 32.098 464 8132.0984648132.098\,464\,8132.098 464 81 2.746 023 6482.7460236482.746\,023\,6482.746 023 648 33.474 068 5733.4740685733.474\,068\,5733.474 068 57 Ours 2.516 871 5322.5168715322.516\,871\,5322.516 871 532 29.475 340 6529.4753406529.475\,340\,6529.475 340 65 2.695 166 2252.6951662252.695\,166\,2252.695 166 225 33.046 739 9333.0467399333.046\,739\,9333.046 739 93 indoor3 outdoor1 EPE $\downarrow$ %Out $\downarrow$ EPE $\downarrow$ %Out $\downarrow$ MultiCM [47] 2.731 729 672.731729672.731\,729\,672.731 729 67 33.218 937 4433.2189374433.218\,937\,4433.218 937 44 1.960 064 5411.9600645411.960\,064\,5411.960 064 541 18.159 503 3618.1595033618.159\,503\,3618.159 503 36 Ours 2.653 875 2772.6538752772.653\,875\,2772.653 875 277 32.241 488 8832.2414888832.241\,488\,8832.241 488 88 1.864 629 221.864629221.864\,629\,221.864 629 22 17.554 100 1417.5541001417.554\,100\,1417.554 100 14
Table 4: Optical flow estimation results on MVSEC [61] ($dt=4$ frames, i.e., 89 ms) with BA noise at $5$ Hz.
4.4 Application to Intensity Reconstruction
In real-world use cases, denoising works as preprocessing steps for other tasks (e.g., deblur, 3D reconstruction, motion estimation, etc.), hence it is important for denoising methods to demonstrate applicability. Although the proposed approach intrinsically improves motion estimation (Sec. 4.3), we further evaluate the denoising outcome via intensity reconstruction, a well-known task in event-based vision, using the E-MLB dataset [10]. Figure 7 collects the results. We use two state-of-the-art approaches for reconstruction: E2VID [45], which does not rely on motion, and EVILIP [58], which utilizes optical flow as prior.
Our method provides flow (Fig. 7, column (b)) and signal events (sharp IWEs in column (c)). Comparing the reconstruction with our event denoising method (column (d)) and the reconstruction from the raw event data (column (e)), the images after event denoising have fewer artifacts than those without denoising, for both reconstruction methods tested. We observe that E2VID is more robust against noise than EVILIP, possibly due to noise events in its training (simulation) data. Due to the ND filters, the quality of the reference frames (last column) is limited, and hence, we discuss the results qualitatively. Nevertheless comparing the reconstructed images to the reference frames, evidences the HDR advantages of event cameras over frame-based ones.
$\tau=0.9$
Refer to caption Refer to caption Refer to caption
$\tau=0.7$
Refer to caption Refer to caption Refer to caption
$\tau=0.5$
Refer to caption Refer to caption Refer to caption
(a) Signal (b) Noise (c) Reconstruction
Figure 8: Sensitivity analysis: effect of different target ratio parameters $\tau$ on image reconstruction using EVILIP [58].
4.5 Sensitivity
Effect of noise ratio. As Sec. 4.2 already showed the sensitivity for denoising accuracy, let us further discuss the sensitivity using image reconstruction. Figure 8 presents the result of the Bicycle sequence with ND64 filter [10]. While the GT image is highly underexposed and not informative (as those on the last column of Fig. 7), the reconstructed images clearly reflect the choice of the signal target ratio $\tau$. As more events are removed, ($i$) the reconstructed image becomes more homogeneous but preserves strong edges, and ($ii$) the removed events (2nd col.) contain more scene edges (i.e., signal). We quantitatively analyze the former in the supplementary using non-reference image quality metrics.
Effect of the kernel size. Figure 9 shows the sensitivity analysis for different kernel sizes. By increasing the size of the Gaussian kernel in the IWE, our method can emphasize contour edges over isolated points, thus better preserving signal events in regions with low IWE intensities (center of the image, corresponding to far away scene points).
Refer to caption Refer to caption Refer to caption
(a) Ref. events (b) $k=1$ px (c) $k=3$ px
Figure 9: Results of different Gaussian kernel sizes $k$ for sensitivity on different depth scenes.
4.6 Ablation of Motion Estimator
The proposed approach can be combined with other estimation methods, such as deep-neural–network (DNN) flow estimators. To this end, we show an example of E-RAFT [21]. The model-based method alone [47] is known to output a degraded flow in night sequences with lots of noise, such as zurich_city_12a [20]. As shown in Fig. 10, our method can identify and discard noisy events. E-RAFT provides sharper and more natural edges in IWEs after denoising compared to MultiCM [47] (thanks to its more accurate flow). ESR scores are similar: $2.39$ for E-RAFT and $2.45$ for MultiCM. While we focus on the inference of DNN in the proposed pipeline, future research could look into accommodating training within the proposed joint estimation framework.
4.7 Runtime
Using $30$k events from the ECD dataset ($240\times 180$ px) on a Mac M1 CPU (2020), the denoising operations (scoring and ranking) take less than $0.1$ s extra computation in total, out of 2 s of the entire processing time (i.e., $5$% computation increase with respect to the base method). Nonetheless, accommodating learning-based approaches in our framework could be key to the trade-off between accuracy and speed.
Refer to caption Refer to caption
Raw Event IWE (No Warp)
MultiCM [47]
Refer to caption Refer to caption Refer to caption
E-RAFT [21]
Refer to caption Refer to caption Refer to caption
(a) Signal Events (b) IWE (b) Flow Estimation
Figure 10: Effect of adopting two different motion estimators (E-RAFT or MultiCM) on the DSEC night scene (zurich_12a) [20].
5 Conclusion
Motion is an intrinsic part of event camera data, since scene edges cannot be sensed without it. We propose the first joint estimation framework for denoising and generic motion estimation, which connects two topics that have traditionally been disconnected in event-based vision. The experiments show that the proposed method achieves state-of-the-art results on the E-MLB benchmark and competitive results on the DND21 benchmark for denoising, while showing its efficacy on motion estimation and intensity reconstruction tasks. Moreover, the proposed method is flexible, as it allows replacing the 1-step motion estimation of CMax with any other motion estimator, such as DNNs. We believe that the proposed approach contributes not only to strengthening the theory of event-data denoising but also to expanding practical denoising use-cases via the code published.
Acknowledgments
We would like to thank Dr. B. Jiang for useful discussions. This research was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2002/1 “Science of Intelligence” – project number 390523135.
6 Supplementary
6.1 Video
We encourage readers to inspect the attached video, which summarizes the method and the results.
$\text{iter}=0$
Refer to caption Refer to caption Refer to caption
$\text{iter}=20$
Refer to caption Refer to caption Refer to caption
$\text{iter}=100$
Refer to caption Refer to caption Refer to caption
(a) Signal (b) Noise (c) IWE
Figure 11: Evolutions of signal, noise, and motion (IWE) during optimization. The edge structure (e.g., green boxes in (b) Noise) converges to move to signal events, while the CMax converges to the sharp IWE (i.e., expected motion parameters).
(a) Signal (first iter)
(b) Signal (last iter)
(c) IWE (last iter)
(d) Evolution of denoising metrics
(e) Evolution of loss function
Figure 12: Results on HD real-world data and intermediate denoising values (TPR, FPR and ESR metrics) during optimization.
1Hz 3Hz 5Hz 7Hz 10Hz hotel driving hotel driving hotel driving hotel driving hotel driving Model-based KNoise [30] $0.6773$ $0.6296$ $0.6521$ $0.6230$ $0.6703$ $0.6235$ $0.6583$ $0.6164$ $0.6413$ $0.6142$ DWF [24] $0.9268$ $0.7409$ $0.8930$ $0.7099$ $0.8620$ $0.6901$ $0.8338$ $0.6747$ $0.7958$ $0.6563$ BAF [9] $0.9535$ $0.8479$ $0.9197$ $0.8155$ $0.8916$ $0.7930$ $0.8662$ $0.7732$ $0.8366$ $0.7479$ Ynoise [14] $0.9690$ $0.9409$ $0.9517$ $0.9240$ $0.9234$ $0.9093$ $0.9177$ $0.8972$ $0.8987$ $0.8800$ TS [33] $0.9716$ $0.9307$ $0.9721$ $0.9260$ $0.9606$ $0.9270$ $0.9654$ $0.9241$ $0.9620$ $0.9202$ Ours 1.013 802 194 317 176 21.01380219431717621.013\,802\,194\,317\,176\,21.013 802 194 317 176 2 0.882 119 693 133 694 80.88211969313369480.882\,119\,693\,133\,694\,80.882 119 693 133 694 8 0.967 937 633 567 141 90.96793763356714190.967\,937\,633\,567\,141\,90.967 937 633 567 141 9 0.850 838 015 623 797 50.85083801562379750.850\,838\,015\,623\,797\,50.850 838 015 623 797 5 0.962 515 815 619 444 40.96251581561944440.962\,515\,815\,619\,444\,40.962 515 815 619 444 4 0.855 398 484 638 116 20.85539848463811620.855\,398\,484\,638\,116\,20.855 398 484 638 116 2 0.951 406 776 235 535 30.95140677623553530.951\,406\,776\,235\,535\,30.951 406 776 235 535 3 0.846 774 906 545 3740.8467749065453740.846\,774\,906\,545\,3740.846 774 906 545 374 0.960 564 724 999 644 20.96056472499964420.960\,564\,724\,999\,644\,20.960 564 724 999 644 2 0.836 273 555 102 2360.8362735551022360.836\,273\,555\,102\,2360.836 273 555 102 236 Learning EDnCNN [4] $0.9573$ $0.8873$ $0.9371$ $0.8771$ $0.9365$ $0.8748$ $0.9254$ $0.8654$ $0.9006$ $0.874$ MLPF [24] $0.9704$ $0.8887$ $0.9718$ $0.8873$ $0.9704$ $0.8845$ $0.9691$ $0.8817$ $0.9634$ $0.8761$ EDformer [28] $0.9928$ $0.9541$ $0.9891$ $0.9472$ $0.9845$ $0.9424$ $0.9792$ $0.9343$ $0.9699$ $0.9264$
[…]
dynamic_rot boxes_rot RMS $\downarrow$ FWL $\uparrow$ RMS $\downarrow$ FWL $\uparrow$ $5$ Hz Noise CMax [17] 20.000 709 6820.0007096820.000\,709\,6820.000 709 68 1.258 796 0971.2587960971.258\,796\,0971.258 796 097 124.641 377124.641377124.641\,377124.641 377 1.128 529 5931.1285295931.128\,529\,5931.128 529 593 – w/ Init. 8.274 876 2458.2748762458.274\,876\,2458.274 876 245 1.273 697 0451.2736970451.273\,697\,0451.273 697 045 20.659 240 3320.6592403320.659\,240\,3320.659 240 33 1.214 159 4591.2141594591.214\,159\,4591.214 159 459 Downsampling $90$% 8.807 703 958.807703958.807\,703\,958.807 703 95 1.272 907 8221.2729078221.272\,907\,8221.272 907 822 124.868 639 7124.8686397124.868\,639\,7124.868 639 7 1.107 758 3951.1077583951.107\,758\,3951.107 758 395 – w/ Init. 8.226 219 4448.2262194448.226\,219\,4448.226 219 444 1.273 725 3071.2737253071.273\,725\,3071.273 725 307 20.618 615 8620.6186158620.618\,615\,8620.618 615 86 1.214 188 4551.2141884551.214\,188\,4551.214 188 455 Downsampling $80$% 9.964 567 2999.9645672999.964\,567\,2999.964 567 299 1.270 984 5391.2709845391.270\,984\,5391.270 984 539 161.810 797 8161.8107978161.810\,797\,8161.810 797 8 1.060 671 $1$.0606711111.060\,671\,$1$.060 671 111 – w/ Init. 8.243 565 3688.2435653688.243\,565\,3688.243 565 368 1.273 741 8421.2737418421.273\,741\,8421.273 741 842 20.797 769 5620.7977695620.797\,769\,5620.797 769 56 1.214 169 4341.2141694341.214\,169\,4341.214 169 434 Downsampling $70$% 14.398 878 9114.3988789114.398\,878\,9114.398 878 91 1.262 100 3441.2621003441.262\,100\,3441.262 100 344 184.511 383 4184.5113834184.511\,383\,4184.511 383 4 1.030 485 5451.0304855451.030\,485\,5451.030 485 545 – w/ Init. 8.231 444 7738.2314447738.231\,444\,7738.231 444 773 1.273 783 9431.2737839431.273\,783\,9431.273 783 943 23.678 913 0923.6789130923.678\,913\,0923.678 913 09 1.210 643 261.210643261.210\,643\,261.210 643 26 Ours $90$% 8.521 574 0738.5215740738.521\,574\,0738.521 574 073 1.273 446 2961.2734462961.273\,446\,2961.273 446 296 97.584 574 8497.5845748497.584\,574\,8497.584 574 84 1.140 880 661.140880661.140\,880\,661.140 880 66 – w/ Init. 8.188 607 2268.1886072268.188\,607\,2268.188 607 226 1.273 725 2991.2737252991.273\,725\,2991.273 725 299 20.604 334 6220.6043346220.604\,334\,6220.604 334 62 1.214 179 0181.2141790181.214\,179\,0181.214 179 018 Ours $80$% 8.511 006 6718.5110066718.511\,006\,6718.511 006 671 1.273 293 7311.2732937311.273\,293\,7311.273 293 731 131.355 614 8131.3556148131.355\,614\,8131.355 614 8 1.102 904 5081.1029045081.102\,904\,5081.102 904 508 – w/ Init. 8.170 039 698.170039698.170\,039\,698.170 039 69 1.273 790 7891.2737907891.273\,790\,7891.273 790 789 20.861 568 0320.8615680320.861\,568\,0320.861 568 03 1.214 094 851.214094851.214\,094\,851.214 094 85 Ours $70$% 9.180 407 8239.1804078239.180\,407\,8239.180 407 823 1.272 074 7881.2720747881.272\,074\,7881.272 074 788 163.676 240 5163.6762405163.676\,240\,5163.676 240 5 1.061 940 9851.0619409851.061\,940\,9851.061 940 985 – w/ Init. 8.086 262 248.086262248.086\,262\,248.086 262 24 1.273 578 8461.2735788461.273\,578\,8461.273 578 846 21.151 118 9621.1511189621.151\,118\,9621.151 118 96 1.213 969 0641.2139690641.213\,969\,0641.213 969 064 BAF 19.674 893 0519.6748930519.674\,893\,0519.674 893 05 1.259 772 491.259772491.259\,772\,491.259 772 49 125.028 145 1125.0281451125.028\,145\,1125.028 145 1 1.127 308 9021.1273089021.127\,308\,9021.127 308 902 – w/ Init. 8.253 078 2028.2530782028.253\,078\,2028.253 078 202 1.273 636 991.273636991.273\,636\,991.273 636 99 19.550 436 619.550436619.550\,436\,619.550 436 6 1.213 505 481.213505481.213\,505\,481.213 505 48 $1$ Hz CMax [17] 19.395 190 2319.3951902319.395\,190\,2319.395 190 23 1.275 866 1481.2758661481.275\,866\,1481.275 866 148 117.439 798 6117.4397986117.439\,798\,6117.439 798 6 1.143 852 7451.1438527451.143\,852\,7451.143 852 745 – w/ Init. 8.254 263 2638.2542632638.254\,263\,2638.254 263 263 1.289 966 6551.2899666551.289\,966\,6551.289 966 655 20.627 507 7120.6275077120.627\,507\,7120.627 507 71 1.223 032 6881.2230326881.223\,032\,6881.223 032 688 Downsampling $90$% 8.676 346 0658.6763460658.676\,346\,0658.676 346 065 1.289 459 4391.2894594391.289\,459\,4391.289 459 439 110.568 366 3110.5683663110.568\,366\,3110.568 366 3 1.130 495 7731.1304957731.130\,495\,7731.130 495 773 – w/ Init. 8.184 187 8618.1841878618.184\,187\,8618.184 187 861 1.289 611 9121.2896119121.289\,611\,9121.289 611 912 20.620 358 2120.6203582120.620\,358\,2120.620 358 21 1.223 050 3321.2230503321.223\,050\,3321.223 050 332 Ours $90$% 8.505 807 8288.5058078288.505\,807\,8288.505 807 828 1.289 787 0891.2897870891.289\,787\,0891.289 787 089 87.774 541 8687.7745418687.774\,541\,8687.774 541 86 1.158 678 1341.1586781341.158\,678\,1341.158 678 134 – w/ Init. 8.177 036 3548.1770363548.177\,036\,3548.177 036 354 1.290 007 1851.2900071851.290\,007\,1851.290 007 185 20.569 191 6120.5691916120.569\,191\,6120.569 191 61 1.223 052 861.223052861.223\,052\,861.223 052 86 BAF 19.568 905 3719.5689053719.568\,905\,3719.568 905 37 1.275 738 1221.2757381221.275\,738\,1221.275 738 122 117.554 331 4117.5543314117.554\,331\,4117.554 331 4 1.143 238 0781.1432380781.143\,238\,0781.143 238 078 – w/ Init. 8.189 429 3938.1894293938.189\,429\,3938.189 429 393 1.289 532 0351.2895320351.289\,532\,0351.289 532 035 20.713 437 520.713437520.713\,437\,520.713 437 5 1.223 006 3741.2230063741.223\,006\,3741.223 006 374
Table 6: Angular velocity estimation on ECD dataset [39].
6.2 Denoising Convergence During Optimization
To further validate the proposed joint estimation approach, we analyze the convergence during the joint estimation using the ECD dataset in Fig. 11 (see also results in Fig. 5). At the first iteration (i.e., initialization), signal and noise events are classified randomly. As the optimization proceeds, signal events evolve to keep edge structures in the scene, while noise events evolve to drop such edge structures (see the second and third rows). Also, the IWEs converge to sharp edges with correct motion parameters. This example further confirms the efficacy of the joint estimation.
The optimization process on an HD-resolution ($1280\times 720$ px) real-world dataset, TUM-VIE [32], is shown in Fig. 12. The intermediate progress (d)–(e) demonstrates how motion estimation converges and the denoising performance improves, simultaneously.
6.3 Full Results on Denoising DND21 data
Table 5 is the full version of Tab. 2 in the main paper (including added noise at rates of 3 and 7 Hz).
6.4 Full Results on Angular Velocity Estimation
While the quantitative evaluation on angular velocity estimation is summarized in Fig. 6, here, we report the detailed results with different target ratio parameters, also compared with other baselines such as BA Filter [9]. The original CMax degrades due to noise, as reported in previous work (e.g., [3]). The S/N target ratio $\tau$ affects the accuracy: when it is close to the actual value of noise injection, the result of the proposed method becomes better. The amount of artificial noise injected is around $15$ for $5$ Hz and $3$ for $1$ Hz conditions. Although we cannot know the “true” noise level due to the original noise in the ECD sequences, our method constantly produces better accuracy and FWL than other baselines. Please refer to Sec. 4.3 for more discussions about dependency on initialization and comparison with other baselines. The AUCs for the conditions that we test ($\tau=\{0.9,\ldots,0.7\}$) are $0.70$ (“Ours”) and $0.67$ (“Downsampling”).
6.5 Quantitative Evaluation of Intensity Reconstruction
In Secs. 4.4 and 4.5 we show qualitative results of the intensity reconstruction application. Here, we discuss possible quantitative evaluation. The challenge of the quantitative evaluation lies in the quality of reference frames (i.e., “GT”) in the existing dataset as shown in Figs. 2 and 7: the frames become underexposed or blurry due to their limited dynamic range, when event data suffer from more BA noise (i.e., in dark scenes).
Nonetheless, we report non-reference image quality indices for different S/N ratios ($\tau$). Figure 13 reports the scores of Blind/Referenceless Image Spatial Quality Evaluator (BRISQUE) [37] and Naturalness Image Quality Evaluator (NIQE) [38], using Bicycle-ND64-2 sequence (same as Fig. 8). These scores indicate the perceptual quality of images, and smaller is better. Although BRISQUE monotonically increases as the target ratio decreases (i.e., more events are removed), NIQE scores the lowest at $\tau=0.9$, indicating the best quality of the reconstructed image. Although the results potentially suggest that it could estimate the “true” noise ratio in the data using the non-reference indices, which is useful for image reconstruction applications, we leave further evaluation and discussion as future work.
Figure 13: Results of the non-reference image quality indices for image reconstruction. See Fig. 8 for the details.
7 Limitations
Our method expects as parameter a target noise level, but the intrinsic (“true”) noise level depends on the camera, scene, etc. and is therefore unknown. More research in this direction is needed for automatic parameter tuning.
Also, notice that there might be a potential limitation of the denoising metric ESR, since it depends on the motion parameter (see Suppl. Mat. of [28]). Although our denoising efficacy is confirmed from multiple perspectives, we believe that advancing denoising metrics for real-world data (where it gets interesting, as there are no GT labels) is a central direction for future work in event-based vision.

### v2

@@
 Abstract
-Event cameras are emerging vision sensors, whose noise is challenging to characterize.
-Existing denoising methods for event cameras consider other tasks such as motion estimation separately (i.e., sequentially after denoising).
+Event cameras are emerging vision sensors whose noise is challenging to characterize.
+Existing denoising methods for event cameras are often designed in isolation and thus consider other tasks, such as motion estimation, separately (i.e., sequentially after denoising).
 However, motion is an intrinsic part of event data, since scene edges cannot be sensed without motion.
-This work proposes, to the best of our knowledge, the first method that simultaneously estimates motion in its various forms (e.g., ego-motion, optical flow) and noise.
-The method is flexible, as it allows replacing the 1-step motion estimation of the widely-used Contrast Maximization framework with any other motion estimator, such as deep neural networks.
-The experiments show that the proposed method achieves state-of-the-art results on the E-MLB denoising benchmark and competitive results on the DND21 benchmark, while showing its efficacy on motion estimation and intensity reconstruction tasks.
-We believe that the proposed approach contributes to strengthening the theory of event-data denoising, as well as impacting practical denoising use-cases, as we release the code upon acceptance.
+We propose, to the best of our knowledge, the first method that simultaneously estimates motion in its various forms (e.g., ego-motion, optical flow) and noise.
+The method is flexible, as it allows replacing the one-step motion estimation of the widely-used Contrast Maximization framework with any other motion estimator, such as deep neural networks.
+The experiments show that the proposed method achieves state-of-the-art results on the E-MLB denoising benchmark and competitive results on the DND21 benchmark, while demonstrating effectiveness across motion estimation and intensity reconstruction tasks.
+Our approach advances event-data denoising theory and expands practical denoising use-cases via open-source code.
 Project page: https://github.com/tub-rip/ESMD
+{textblock}
+11(2.5, 0.6)
+This paper has been accepted for publication at the
+IEEE International Conference on Computer Vision (ICCV), Honolulu, 2025. ©IEEE
 1 Introduction
-Event cameras are emerging vision sensors that overcome challenges of conventional cameras (e.g., motion blur, limited dynamic range, data and power efficiency), yet they suffer from a considerable amount of noise due to their novelty and operation in low-power (transistor subthreshold) conditions [19].
+Event cameras are emerging vision sensors that overcome shortcomings of conventional cameras (e.g., motion blur, limited dynamic range, data and power efficiency), yet they suffer from a considerable amount of noise due to their novelty and operation in low-power (transistor subthreshold) conditions [19, 36].
 Since event cameras are suitable for many computer vision tasks, especially motion-related tasks, it is paramount to classify event data that is related to motion (i.e., signal) and that is not (i.e., noise).
-However, denoising event data is a challenging problem, since noise properties are not fully characterized to date, and it is not feasible to define ground-truth (GT) noise labels in real-world data recordings.
-Previous work either ($i$) leveraged simulation to generate pure signal (events from edge motion under constant illumination) and noise, or ($ii$) prepared real event data that was expected to be clean, by aggressively removing noise-like events with existing filters and using bright illumination that minimized noise occurrence.
+However, denoising event data is a challenging problem because noise properties have yet to be fully characterized [24, 22] and it is infeasible to define ground-truth (GT) noise labels in real-world data recordings.
+Previous works either leveraged simulation or prepared real event data and aggressively suppressed noise occurrence.
 Figure 1: Overview of the proposed method. We use only raw events as input, and estimate signal events, noise events, and motion (e.g., dense optical flow and ego-motion estimation).
@@
 In this work, we focus on noise due to the leakage inside the camera hardware, such as background activity (BA) events [36].
-Events due to flickering or active lights can be treated as “noise” for methods that consider passive sensing, which are a large majority in computer vision, and are therefore not our focus.
+Events due to flickering or active lights are therefore not our priority.
 Classical denoising works focus on the spatio-temporal correlation of events, as BA events have fewer neighboring events in space-time coordinates than signal events.
@@
 While these approaches are effective for benchmarking, the challenge of obtaining GT labels for practical real-world datasets remains unsolved.
-Our work builds upon a series of successful methods on motion estimation and their applications.
-Event-based motion estimation is a paramount topic since event cameras naturally respond to motion in the scenes [19].
-In particular, Contrast Maximization (CMax) [17, 49] is a state-of-the-art framework for motion estimation, which has various applications, such as rotational motion [15, 23, 44, 41, 31], optical flow [62, 43, 47, 27], intensity reconstruction [58], and SLAM [25, 22, 54].
+Our work builds upon a series of successful methods for motion estimation and their applications.
+Event-based motion estimation is an important topic since event cameras naturally respond to motion in the scenes [19].
+In particular, Contrast Maximization (CMax) [17, 50] is a state-of-the-art framework for motion estimation, which has various applications, such as rotational motion estimation [15, 23, 45, 42, 32], optical flow estimation [64, 44, 48, 28, 27], and motion segmentation [51, 60, 57].
 We investigate the connection and potential benefits of applying the CMax framework to denoising during the estimation process (i.e., simultaneously), demonstrating motion estimation improvements as well as competitive denoising accuracy.
-A prior work, “ProgressiveMotionSeg” [8], has extended the CMax-based segmentation method [50] with extra weights that accounted for how much each event contributed to estimated motion clusters.
-However, [8] mainly focuses on the task of motion segmentation using low-DOF warps (e.g., 2-DOF feature flow) and does not discuss denoising efficacy or evaluate on standard denoising benchmarks.
-In contrast, our method supports various types of motion estimation (both low-DOF and high-DOF warps like optical flow), and considers that noise shall not be used to estimate motion, as from first principles we deem noise and motion to be uncorrelated.
-We propose signal-noise classifiers and show ranking invariance, which results in state-of-the-art denoising accuracy and improvements on motion estimation from noisy data.
+A prior work, “ProgressiveMotionSeg” [8], has extended the CMax-based segmentation method [51] with extra weights that accounted for how much each event contributed to the estimated motion clusters.
+However, [8] mainly focuses on the task of motion segmentation using low degrees-of-freedom (DOF) warps (e.g., 2-DOF feature flow) and does not discuss denoising efficacy or evaluate on standard denoising benchmarks.
+In contrast, our method supports various types of motion estimation (both low-DOF and high-DOF warps, such as optical flow), and considers that noise shall not be used to estimate motion, as from first principles we deem noise and motion to be uncorrelated.
 3 Methodology
-The following sections present the CMax framework (Sec. 3.1), the proposed signal/noise classifier (Sec. 3.2), and the simultaneous denoiser and motion estimator (Sec. 3.3).
 Event cameras consist of independent pixels that generate asynchronous “events” when the logarithmic brightness at the pixel increases or decreases by a predefined contrast sensitivity.
@@
 The CMax framework [17] assumes events $\mathcal{E}\doteq\{e_{k}\}_{k=1}^{N_{e}}$ are caused by moving edges, and it transforms their coordinates according to a motion model $\mathbf{W}$, producing a set of warped events $\mathcal{E}^{\prime}_{t_{\text{ref}}}\doteq\{e^{\prime}_{k}\}_{k=1}^{N_{e}}$ at a reference time $t_{\text{ref}}$:
-$\vspace{-0.5ex}e_{k}\doteq(\mathbf{x}_{k},t_{k},p_{k})\;\,\mapsto\;\,e^{\prime}_{k}\doteq(\mathbf{x}^{\prime}_{k},t_{\text{ref}},p_{k}).\vspace{-0.5ex}$ (1)
+$\vskip-2.58334pte_{k}\doteq(\mathbf{x}_{k},t_{k},p_{k})\;\,\mapsto\;\,e^{\prime}_{k}\doteq(\mathbf{x}^{\prime}_{k},t_{\text{ref}},p_{k}).\vskip-0.86108pt$ (1)
 The warp $\mathbf{x}^{\prime}_{k}=\mathbf{W}(\mathbf{x}_{k},t_{k};\boldsymbol{\theta})$ transports each event from $t_{k}$ to $t_{\text{ref}}$ along the motion curve that passes through it.
 Then, they are aggregated on an image of warped events (IWE):
-$\vspace{-0.5ex}\textstyle I(\mathbf{x};\mathcal{E}^{\prime}_{t_{\text{ref}}},\boldsymbol{\theta})\doteq\sum_{k=1}^{N_{e}}\delta(\mathbf{x}-\mathbf{x}^{\prime}_{k}),\vspace{-0.5ex}$ (2)
+$\vskip-2.58334pt\textstyle I(\mathbf{x};\mathcal{E}^{\prime}_{t_{\text{ref}}},\boldsymbol{\theta})\doteq\sum_{k=1}^{N_{e}}\delta(\mathbf{x}-\mathbf{x}^{\prime}_{k}),\vskip-0.86108pt$ (2)
 where each pixel $\mathbf{x}$ sums the number of warped events $\mathbf{x}^{\prime}_{k}$ that fall within it.
 The Dirac delta is approximated by a Gaussian, $\delta(\mathbf{x}-\boldsymbol{\mu})\approx\mathcal{N}(\mathbf{x};\boldsymbol{\mu},\epsilon^{2}\mathtt{Id})$ with $\epsilon=1$ pixel.
-GasTank (E-MLB) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+GasTank (E-MLB)
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
 LEGO (E-MLB)
@@
 (a) Input (Raw Events) (b) IWE (No Warp) (c) Output (Signal Events) (d) IWE (Signal Events) (e) Estimated Flow (f) Reference Frame
-Figure 2: Qualitative result of denoising and flow estimation on E-MLB and DND21 data.
-The driving sequence has $5$-Hz artificial noise.
-(b) is an alternative visualization of (a) with the same grayscale coding as (d) for ease of comparison.
-Flow color is given in Fig.
+Figure 2: Qualitative result of denoising and flow estimation on E-MLB and DND21.
+The driving sequence has 555-Hz artificial noise.
+(b) is an alternative visualization of (a) with the same grayscale coding as (d).
+Flow is visualized on all events and flow color is given in Fig.
 1.
@@
 Without labels or additional information, events rely on each other to realize this classification.
-To this end, we compute a score $c_{k}$ for each event, rank the events according to their score, and threshold at some value $T$.
+To this end, we compute a score $c_{k}$ for each event, rank the events according to their scores, and threshold at some value $T$.
 The threshold is computed by assuming the amount or ratio of noisy events is known $\eta$, i.e., $T\equiv T(\eta)$, which allows us to have high-level control over the denoising process.
@@
 $\begin{split}\mathcal{E}_{s}\equiv\mathcal{E}_{\text{signal}}&\doteq\{e_{k}\in\mathcal{E}\;|\;c_{k}>T(\eta)\},\\ \mathcal{E}_{n}\equiv\mathcal{E}_{\text{noise}}&\doteq\mathcal{E}\backslash\mathcal{E}_{\text{signal}}\quad(\text{complement of }\mathcal{E}_{\text{signal}}).\end{split}$ (3)
-We studied different choices for the per-event score $c_{k}$ based on the agreement among warped events $\mathcal{E}^{\prime}=\mathcal{E}^{\prime}_{s}\cup\mathcal{E}^{\prime}_{n}$, through their IWEs, and concluded on two.
+We studied different choices for the per-event score $c_{k}$ based on the agreement among warped events $\mathcal{E}^{\prime}=\mathcal{E}^{\prime}_{s}\cup\mathcal{E}^{\prime}_{n}$, through their IWEs, and concluded on two, as follows.
 Letting $I^{s}_{k}\equiv I^{s}_{k}(\mathbf{x}^{\prime}_{k}),I^{n}_{k}\equiv I^{n}_{k}(\mathbf{x}^{\prime}_{k})$ and $I_{k}\equiv I(\mathbf{x}^{\prime}_{k})=I^{s}_{k}+I^{n}_{k}$ be the IWE values of previously classified signal events, noise events and all events, respectively, the two choices are: ($i$) ${}^{1}c_{k}\doteq I_{k}$, and ($ii$) the ratio of signal events ${}^{2}c_{k}\doteq I^{s}_{k}/I_{k}$ (probabilistic viewpoint).
@@
 Classifying events into signal and noise requires knowledge of the true motion, and estimation of the true motion (e.g., using CMax) requires knowledge of the signal events because, by definition, noise events carry no information about motion.
-Hence, it is a chicken-and-egg problem.
+Hence, it is a circular dependency problem.
 We approach its solution iteratively, as shown in Fig.
@@
 4.6).
-E-MLB (Day) E-MLB (Night) DND21 ND1 ND4 ND16 ND64 ND1 ND4 ND16 ND64 Model-based Raw $0.821$ $0.824$ $0.815$ $0.786$ $0.89$ $0.824$ $0.786$ $0.768$ $0.869$ BAF [9] $0.861$ $0.869$ $0.876$ $0.89$ $0.946$ $0.973$ $0.992$ $0.942$ $0.92$ KNoise [30] $0.846$ $0.837$ $0.83$ $0.807$ $0.954$ $0.956$ $0.871$ $0.817$ $0.887$ DWF [24] $0.878$ $0.876$ $0.866$ $0.865$ $0.923$ $0.962$ $0.988$ $0.932$ $0.905$ EvFlow [53] $0.848$ $0.878$ $0.868$ $0.833$ $0.969$ $0.983$ $0.889$ $0.797$ $1.006$ Ynoise [14] $0.866$ $0.863$ $0.857$ $0.821$ $1.009$ $0.943$ $0.875$ $0.792$ $0.966$ TS [33] $0.877$ $0.887$ $0.87$ $0.837$ $1.033$ $0.944$ $0.886$ $0.797$ $0.985$ IETS [5] $0.772$ $0.785$ $0.777$ $0.753$ $0.950$ $0.823$ $0.804$ $0.711$ $0.900$ GEF [11] $1.051$ $0.938$ $0.935$ $0.927$ $1.027$ $0.955$ $0.946$ $0.935$ $0.932$ Ours 0.938 254 384 292 974 70.93825438429297470.938\,254\,384\,292\,974\,70.938 254 384 292 974 7 0.958 464 335 277 024 80.95846433527702480.958\,464\,335\,277\,024\,80.958 464 335 277 024 8 0.985 896 450 122 991 80.98589645012299180.985\,896\,450\,122\,991\,80.985 896 450 122 991 8 0.950 081 880 240 459 60.95008188024045960.950\,081\,880\,240\,459\,60.950 081 880 240 459 6 1.036 831 015 641 0911.0368310156410911.036\,831\,015\,641\,0911.036 831 015 641 091 0.961 426 679 598 1070.9614266795981070.961\,426\,679\,598\,1070.961 426 679 598 107 0.944 527 953 022 538 40.94452795302253840.944\,527\,953\,022\,538\,40.944 527 953 022 538 4 0.932 362 037 844 1260.9323620378441260.932\,362\,037\,844\,1260.932 362 037 844 126 0.992 069 121 10.99206912110.992\,069\,121\,10.992 069 121 1 Learning EventZoom [12] $0.996$ $0.988$ $0.996$ $0.97$ $1.055$ $1.007$ $1.01$ $0.988$ $1.059$ EDnCNN [4] $0.887$ $0.908$ $0.903$ $0.912$ $1.001$ $1.024$ $1.079$ $1.086$ $0.977$ MLPF [24] $0.851$ $0.855$ $0.846$ $0.84$ $0.926$ $0.928$ $0.91$ $0.906$ $0.944$ EDformer [28] $0.952$ $0.955$ $0.956$ $0.942$ $1.048$ $1.019$ $1.076$ $1.099$ $1.041$
-[…]
-1Hz 5Hz 10Hz hotel driving hotel driving hotel driving Model-based KNoise [30] $0.6773$ $0.6296$ $0.6703$ $0.6235$ $0.6413$ $0.6142$ DWF [24] $0.9268$ $0.7409$ $0.8620$ $0.6901$ $0.7958$ $0.6563$ BAF [9] $0.9535$ $0.8479$ $0.8916$ $0.7930$ $0.8366$ $0.7479$ Ynoise [14] $0.9690$ $0.9409$ $0.9234$ $0.9093$ $0.8987$ $0.8800$ TS [33] $0.9716$ $0.9307$ $0.9606$ $0.9270$ $0.9620$ $0.9202$ Ours 1.013 802 194 317 176 21.01380219431717621.013\,802\,194\,317\,176\,21.013 802 194 317 176 2 0.882 119 693 133 694 80.88211969313369480.882\,119\,693\,133\,694\,80.882 119 693 133 694 8 0.962 515 815 619 444 40.96251581561944440.962\,515\,815\,619\,444\,40.962 515 815 619 444 4 0.855 398 484 638 116 20.85539848463811620.855\,398\,484\,638\,116\,20.855 398 484 638 116 2 0.960 564 724 999 644 20.96056472499964420.960\,564\,724\,999\,644\,20.960 564 724 999 644 2 0.836 273 555 102 2360.8362735551022360.836\,273\,555\,102\,2360.836 273 555 102 236 Learning EDnCNN [4] $0.9573$ $0.8873$ $0.9365$ $0.8748$ $0.9006$ $0.874$ MLPF [24] $0.9704$ $0.8887$ $0.9704$ $0.8845$ $0.9634$ $0.8761$ EDformer [28] $0.9928$ $0.9541$ $0.9845$ $0.9424$ $0.9699$ $0.9264$
-[…]
-Figure 3: ROC curves on the DND 21 dataset sequences. Dash lines represent learning-based methods (see also Tab. 2).
-Figure 4: Comparison of various denoising methods on the two DND21 sequences (driving and hotel) at two different noise levels.
 4.1 Datasets, Metrics, and Hyper-parameters
@@
 E-MLB [10] is a large-scale, de-facto dataset to benchmark denoising.
-It consists of $100$ sequences with 4 different brightness levels using neutral-density (ND) filters on a DAVIS346 camera ($346\times 260$ px) during day and night.
+It consists of 100100100 sequences with 4 different brightness levels using neutral-density (ND) filters on a DAVIS346 camera (346×260346$\times$ 260346 × 260 px) during day and night.
 Historically, the two sequences (hotel and driving) of the DND21 dataset [24] have been widely used for denoising evaluation [24, 10, 29].
@@
 Combined with other pure BA noise sequences, the dataset has event-wise noise annotations that are useful for evaluation.
-We use the above sequences with different noise rates, $1,3,5,7,10$ Hz per pixel, following prior work [28].
+We use the above sequences with different noise rates, 1,3,5,7,101,3,5,7,101 , 3 , 5 , 7 , 10 Hz per pixel, following prior work [29].
+E-MLB (Day) E-MLB (Night) DND21 ND1 ND4 ND16 ND64 ND1 ND4 ND16 ND64 Model-based Raw 0.8210.8210.821 0.8240.8240.824 0.8150.8150.815 0.7860.7860.786 0.890.890.89 0.8240.8240.824 0.7860.7860.786 0.7680.7680.768 0.8690.8690.869 BAF [9] 0.8610.8610.861 0.8690.8690.869 0.8760.8760.876 0.890.890.89 0.9460.9460.946 0.9730.9730.973 0.9920.9920.992 0.9420.9420.942 0.920.920.92 TS [34] 0.8770.8770.877 0.8870.8870.887 0.870.870.87 0.8370.8370.837 1.0331.0331.033 0.9440.9440.944 0.8860.8860.886 0.7970.7970.797 0.9850.9850.985 KNoise [31] 0.8460.8460.846 0.8370.8370.837 0.830.830.83 0.8070.8070.807 0.9540.9540.954 0.9560.9560.956 0.8710.8710.871 0.8170.8170.817 0.8870.8870.887 EvFlow [54] 0.8480.8480.848 0.8780.8780.878 0.8680.8680.868 0.8330.8330.833 0.9690.9690.969 0.9830.9830.983 0.8890.8890.889 0.7970.7970.797 1.0061.0061.006 IETS [5] 0.7720.7720.772 0.7850.7850.785 0.7770.7770.777 0.7530.7530.753 0.9500.9500.950 0.8230.8230.823 0.8040.8040.804 0.7110.7110.711 0.9000.9000.900 Ynoise [14] 0.8660.8660.866 0.8630.8630.863 0.8570.8570.857 0.8210.8210.821 1.0091.0091.009 0.9430.9430.943 0.8750.8750.875 0.7920.7920.792 0.9660.9660.966 GEF [11] 1.0511.0511.051 0.9380.9380.938 0.9350.9350.935 0.9270.9270.927 1.0271.0271.027 0.9550.9550.955 0.9460.9460.946 0.9350.9350.935 0.9320.9320.932 DWF [24] 0.8780.8780.878 0.8760.8760.876 0.8660.8660.866 0.8650.8650.865 0.9230.9230.923 0.9620.9620.962 0.9880.9880.988 0.9320.9320.932 0.9050.9050.905 Ours 0.938 254 384 292 974 70.938\,254\,384\,292\,974\,70.938 254 384 292 974 7 0.958 464 335 277 024 80.958\,464\,335\,277\,024\,80.958 464 335 277 024 8 0.985 896 450 122 991 80.985\,896\,450\,122\,991\,80.985 896 450 122 991 8 0.950 081 880 240 459 60.950\,081\,880\,240\,459\,60.950 081 880 240 459 6 1.036 831 015 641 0911.036\,831\,015\,641\,0911.036 831 015 641 091 0.961 426 679 598 1070.961\,426\,679\,598\,1070.961 426 679 598 107 0.944 527 953 022 538 40.944\,527\,953\,022\,538\,40.944 527 953 022 538 4 0.932 362 037 844 1260.932\,362\,037\,844\,1260.932 362 037 844 126 0.992 069 121 10.992\,069\,121\,10.992 069 121 1 Learning EDnCNN [4] 0.8870.8870.887 0.9080.9080.908 0.9030.9030.903 0.9120.9120.912 1.0011.0011.001 1.0241.0241.024 1.0791.0791.079 1.0861.0861.086 0.9770.9770.977 EventZoom [12] 0.9960.9960.996 0.9880.9880.988 0.9960.9960.996 0.970.970.97 1.0551.0551.055 1.0071.0071.007 1.011.011.01 0.9880.9880.988 1.0591.0591.059 MLPF [24] 0.8510.8510.851 0.8550.8550.855 0.8460.8460.846 0.840.840.84 0.9260.9260.926 0.9280.9280.928 0.910.910.91 0.9060.9060.906 0.9440.9440.944 EDformer [29] 0.9520.9520.952 0.9550.9550.955 0.9560.9560.956 0.9420.9420.942 1.0481.0481.048 1.0191.0191.019 1.0761.0761.076 1.0991.0991.099 1.0411.0411.041
+[…]
+1Hz 5Hz 10Hz hotel driving hotel driving hotel driving Model-based BAF [9] 0.95350.95350.9535 0.84790.84790.8479 0.89160.89160.8916 0.79300.79300.7930 0.83660.83660.8366 0.74790.74790.7479 TS [34] 0.97160.97160.9716 0.93070.93070.9307 0.96060.96060.9606 0.92700.92700.9270 0.96200.96200.9620 0.92020.92020.9202 KNoise [31] 0.67730.67730.6773 0.62960.62960.6296 0.67030.67030.6703 0.62350.62350.6235 0.64130.64130.6413 0.61420.61420.6142 Ynoise [14] 0.96900.96900.9690 0.94090.94090.9409 0.92340.92340.9234 0.90930.90930.9093 0.89870.89870.8987 0.88000.88000.8800 DWF [24] 0.92680.92680.9268 0.74090.74090.7409 0.86200.86200.8620 0.69010.69010.6901 0.79580.79580.7958 0.65630.65630.6563 Ours 1.013 802 194 317 176 21.013\,802\,194\,317\,176\,21.013 802 194 317 176 2 0.882 119 693 133 694 80.882\,119\,693\,133\,694\,80.882 119 693 133 694 8 0.962 515 815 619 444 40.962\,515\,815\,619\,444\,40.962 515 815 619 444 4 0.855 398 484 638 116 20.855\,398\,484\,638\,116\,20.855 398 484 638 116 2 0.960 564 724 999 644 20.960\,564\,724\,999\,644\,20.960 564 724 999 644 2 0.836 273 555 102 2360.836\,273\,555\,102\,2360.836 273 555 102 236 Learning EDnCNN [4] 0.95730.95730.9573 0.88730.88730.8873 0.93650.93650.9365 0.87480.87480.8748 0.90060.90060.9006 0.8740.8740.874 MLPF [24] 0.97040.97040.9704 0.88870.88870.8887 0.97040.97040.9704 0.88450.88450.8845 0.96340.96340.9634 0.87610.87610.8761 EDformer [29] 0.99280.99280.9928 0.95410.95410.9541 0.98450.98450.9845 0.94240.94240.9424 0.96990.96990.9699 0.92640.92640.9264
+[…]
+Figure 3: ROC curves on the DND 21 dataset sequences. Dash lines represent learning-based methods (see also Tab. 2).
+Figure 4: Comparison of denoising methods on the two DND21 sequences at two different noise levels.
 The ECD dataset [40] is a standard dataset for various tasks including camera ego-motion estimation [15, 62, 47, 41, 61, 49].
-Using a DAVIS240C camera ($240\times 180$ px [7]), each sequence provides events, frames, calibration information, IMU data, and ground truth (GT) camera poses (at $200$ Hz).
+Using a DAVIS240C camera (240×180240$\times$ 180240 × 180 px [7]), each sequence provides events, frames, calibration information, IMU data, and ground truth (GT) camera poses (at 200200200 Hz).
 We use boxes_rotation and dynamic_rotation sequences to evaluate rotational motion estimation accuracy.
@@
 Given signal and noise GT labels, the ROC shows the True-Positive Rate and False-Positive Rate over several S/N ratios.
-The AUC of the ROC becomes closer to $1$, the better the denoiser works.
+The AUC of the ROC becomes closer to 111, the better the denoiser works.
 The MESR is a denoising metric that does not require GT labels [10].
@@
 The higher, the better.
-We use the RMS error for angular velocity estimation and the EPE for optical flow estimation, following previous works [41, 23, 25, 49], as well as the FWL metric [51], which is the relative variance (sharpness) of the IWE with respect to that of the identity warp.
-Hyper-parameters.
-For the denoising experiments, we test various values $\tau=\{0.9,\ldots,0.1\}$ to calculate the ROC on DND21 and $\tau=\{0.9,\ldots,0.7\}$ for the RMS on ECD.
-For E-MLB benchmarking, we fix the number of signal events to follow the prior work [10].
-To analyze the proposed pipeline in the CMax framework, we use model-based rotational motion estimation [15] and tile-based optical flow estimation [47] approaches.
-We use the magnitude of the IWE gradient [18] as the CMax objective function.
+Furthermore, we use the RMS error for angular velocity estimation and the EPE for optical flow estimation, following previous works [42, 23, 25, 50], as well as the FWL metric [52], which is the relative variance (sharpness) of the IWE with respect to that of the identity warp.
+Further details about the experimental settings (e.g., hyper-parameters) are in the supplementary.
 4.2 Denoising Results
@@
 Notice that, however, learning-based approaches utilize additional information (GT labels) during training, while our method does not.
-Nonetheless, the results demonstrate the efficacy of our approach, which contributes to the theoretical foundation for the denoising methodology of event data.
+Nonetheless, the results demonstrate the efficacy of our approach, which contributes to the theoretical foundation of event-denoising methodologies.
 We discuss further details of the evaluation metrics (ESR) in Sec.
-7.
+5.
 Qualitative results are shown in Fig.
@@
 Reconstruction on denoised events produces higher quality images (i.e., fewer artifacts) than on raw event data.
-dynamic_rot boxes_rot RMS $\downarrow$ FWL $\uparrow$ RMS $\downarrow$ FWL $\uparrow$ $5$ Hz Noise CMax [17] 20.000 709 6820.0007096820.000\,709\,6820.000 709 68 1.258 796 0971.2587960971.258\,796\,0971.258 796 097 124.641 377124.641377124.641\,377124.641 377 1.128 529 5931.1285295931.128\,529\,5931.128 529 593 – w/ Init.
-8.274 876 2458.2748762458.274\,876\,2458.274 876 245 1.273 697 0451.2736970451.273\,697\,0451.273 697 045 20.659 240 3320.6592403320.659\,240\,3320.659 240 33 1.214 159 4591.2141594591.214\,159\,4591.214 159 459 Downsampling (Best) 8.807 703 958.807703958.807\,703\,958.807 703 95 1.272 907 8221.2729078221.272\,907\,8221.272 907 822 124.868 639 7124.8686397124.868\,639\,7124.868 639 7 1.107 758 3951.1077583951.107\,758\,3951.107 758 395 – w/ Init.
-8.226 219 4448.2262194448.226\,219\,4448.226 219 444 1.273 725 3071.2737253071.273\,725\,3071.273 725 307 20.618 615 8620.6186158620.618\,615\,8620.618 615 86 1.214 188 4551.2141884551.214\,188\,4551.214 188 455 Ours (Best) 8.521 574 0738.5215740738.521\,574\,0738.521 574 073 1.273 446 2961.2734462961.273\,446\,2961.273 446 296 97.584 574 8497.5845748497.584\,574\,8497.584 574 84 1.140 880 661.140880661.140\,880\,661.140 880 66 – w/ Init.
-$8.170$ 1.273 725 2991.2737252991.273\,725\,2991.273 725 299 20.604 334 6220.6043346220.604\,334\,6220.604 334 62 1.214 179 0181.2141790181.214\,179\,0181.214 179 018 BAF 19.674 893 0519.6748930519.674\,893\,0519.674 893 05 1.259 772 491.259772491.259\,772\,491.259 772 49 125.028 145 1125.0281451125.028\,145\,1125.028 145 1 1.127 308 9021.1273089021.127\,308\,9021.127 308 902 – w/ Init.
-8.253 078 2028.2530782028.253\,078\,2028.253 078 202 1.273 636 991.273636991.273\,636\,991.273 636 99 20.550 436 620.550436620.550\,436\,620.550 436 6 1.213 505 481.213505481.213\,505\,481.213 505 48
+dynamic_rot boxes_rot RMS $\downarrow$ FWL $\uparrow$ RMS $\downarrow$ FWL $\uparrow$ 555 Hz Noise CMax [17] 20.000 709 6820.000\,709\,6820.000 709 68 1.258 796 0971.258\,796\,0971.258 796 097 124.641 377124.641\,377124.641 377 1.128 529 5931.128\,529\,5931.128 529 593 – w/ Init.
+8.274 876 2458.274\,876\,2458.274 876 245 1.273 697 0451.273\,697\,0451.273 697 045 20.659 240 3320.659\,240\,3320.659 240 33 1.214 159 4591.214\,159\,4591.214 159 459 Downsampling (Best) 8.807 703 958.807\,703\,958.807 703 95 1.272 907 8221.272\,907\,8221.272 907 822 124.868 639 7124.868\,639\,7124.868 639 7 1.107 758 3951.107\,758\,3951.107 758 395 – w/ Init.
+8.226 219 4448.226\,219\,4448.226 219 444 1.273 725 3071.273\,725\,3071.273 725 307 20.618 615 8620.618\,615\,8620.618 615 86 1.214 188 4551.214\,188\,4551.214 188 455 Ours (Best) 8.521 574 0738.521\,574\,0738.521 574 073 1.273 446 2961.273\,446\,2961.273 446 296 97.584 574 8497.584\,574\,8497.584 574 84 1.140 880 661.140\,880\,661.140 880 66 – w/ Init.
+8.1708.1708.170 1.273 725 2991.273\,725\,2991.273 725 299 20.604 334 6220.604\,334\,6220.604 334 62 1.214 179 0181.214\,179\,0181.214 179 018 BAF 19.674 893 0519.674\,893\,0519.674 893 05 1.259 772 491.259\,772\,491.259 772 49 125.028 145 1125.028\,145\,1125.028 145 1 1.127 308 9021.127\,308\,9021.127 308 902 – w/ Init.
+8.253 078 2028.253\,078\,2028.253 078 202 1.273 636 991.273\,636\,991.273 636 99 20.550 436 620.550\,436\,620.550 436 6 1.213 505 481.213\,505\,481.213 505 48
 Table 3: Angular velocity estimation on ECD dataset [40].
@@
 We used the classical ECD dataset and injected BA noise from the DND21 recordings with different noise rates.
-As no prior work simultaneously estimates motion and noise, we compare against sequential filters: random sampling (of the input data with a uniform distribution) and BAF [9].
+As no prior work simultaneously estimates motion and noise, we compare against sequential filters: random sampling (of the input data with a uniform distribution) and BAF [9] that is the most widely-used denoising method to date.
 As shown in the IWEs, our method produces the sharpest edges (cf.
@@
 4, using the MVSEC dataset [63], with BA noise from the DND21 dataset.
-Due to the $5$Hz BA noise, the baseline method MultiCM [47] degrades.
+Due to the  [^5] Hz BA noise, the baseline method MultiCM [48] degrades.
 Our denoising method improves the flow accuracy, showcasing the efficacy of simultaneous noise estimation.
@@
 Hence, we only evaluate motion estimation metrics, not ROC or AUC.
-indoor1 indoor2 EPE $\downarrow$ %Out $\downarrow$ EPE $\downarrow$ %Out $\downarrow$ MultiCM [47] 2.675 668 1172.6756681172.675\,668\,1172.675 668 117 32.098 464 8132.0984648132.098\,464\,8132.098 464 81 2.746 023 6482.7460236482.746\,023\,6482.746 023 648 33.474 068 5733.4740685733.474\,068\,5733.474 068 57 Ours 2.516 871 5322.5168715322.516\,871\,5322.516 871 532 29.475 340 6529.4753406529.475\,340\,6529.475 340 65 2.695 166 2252.6951662252.695\,166\,2252.695 166 225 33.046 739 9333.0467399333.046\,739\,9333.046 739 93 indoor3 outdoor1 EPE $\downarrow$ %Out $\downarrow$ EPE $\downarrow$ %Out $\downarrow$ MultiCM [47] 2.731 729 672.731729672.731\,729\,672.731 729 67 33.218 937 4433.2189374433.218\,937\,4433.218 937 44 1.960 064 5411.9600645411.960\,064\,5411.960 064 541 18.159 503 3618.1595033618.159\,503\,3618.159 503 36 Ours 2.653 875 2772.6538752772.653\,875\,2772.653 875 277 32.241 488 8832.2414888832.241\,488\,8832.241 488 88 1.864 629 221.864629221.864\,629\,221.864 629 22 17.554 100 1417.5541001417.554\,100\,1417.554 100 14
-Table 4: Optical flow estimation results on MVSEC [61] ($dt=4$ frames, i.e., 89 ms) with BA noise at $5$ Hz.
+indoor1 indoor2 EPE $\downarrow$ %Out $\downarrow$ EPE $\downarrow$ %Out $\downarrow$ MultiCM [48] 2.675 668 1172.675\,668\,1172.675 668 117 32.098 464 8132.098\,464\,8132.098 464 81 2.746 023 6482.746\,023\,6482.746 023 648 33.474 068 5733.474\,068\,5733.474 068 57 Ours 2.516 871 5322.516\,871\,5322.516 871 532 29.475 340 6529.475\,340\,6529.475 340 65 2.695 166 2252.695\,166\,2252.695 166 225 33.046 739 9333.046\,739\,9333.046 739 93 indoor3 outdoor1 EPE $\downarrow$ %Out $\downarrow$ EPE $\downarrow$ %Out $\downarrow$ MultiCM [48] 2.731 729 672.731\,729\,672.731 729 67 33.218 937 4433.218\,937\,4433.218 937 44 1.960 064 5411.960\,064\,5411.960 064 541 18.159 503 3618.159\,503\,3618.159 503 36 Ours 2.653 875 2772.653\,875\,2772.653 875 277 32.241 488 8832.241\,488\,8832.241 488 88 1.864 629 221.864\,629\,221.864 629 22 17.554 100 1417.554\,100\,1417.554 100 14
+Table 4: Optical flow estimation results on MVSEC [63] ($dt=4$ frames, i.e., 89 ms) with BA noise at 555 Hz.
 4.4 Application to Intensity Reconstruction
@@
 E-RAFT provides sharper and more natural edges in IWEs after denoising compared to MultiCM [48] (thanks to its more accurate flow).
-ESR scores are similar: $2.39$ for E-RAFT and $2.45$ for MultiCM.
+ESR scores are similar: 2.392.392.39 for E-RAFT and 2.452.452.45 for MultiCM.
 While we focus on the inference of DNN in the proposed pipeline, future research could look into accommodating training within the proposed joint estimation framework.
 4.7 Runtime
-Using $30$k events from the ECD dataset ($240\times 180$ px) on a Mac M1 CPU (2020), the denoising operations (scoring and ranking) take less than $0.1$ s extra computation in total, out of 2 s of the entire processing time (i.e., $5$% computation increase with respect to the base method).
+Using 303030k events from the ECD dataset (240×180240$\times$ 180240 × 180 px) on a Mac M1 CPU (2020), the denoising operations (scoring and ranking) take less than 0.10.10.1 s extra computation in total, out of 2 s of the entire processing time (i.e., 555% computation increase with respect to the base method).
 Nonetheless, accommodating learning-based approaches in our framework could be key to the trade-off between accuracy and speed.
@@
 Refer to caption Refer to caption Refer to caption
-(a) Signal Events (b) IWE (b) Flow Estimation
+(a) Signal Events (b) IWE (c) Flow Estimation
 Figure 10: Effect of adopting two different motion estimators (E-RAFT or MultiCM) on the DSEC night scene (zurich_12a) [20].
+Despite small IWE differences (e.g., in the car’s hood), the proposed framework supports several motion estimators.
+5 Limitations
+Our method expects as parameter a target noise level, but the intrinsic (“true”) noise level depends on the camera, scene, etc.
+and is therefore unknown.
+More research in this direction is needed for automatic parameter tuning.
+Also, notice that there might be a potential limitation of the denoising metric ESR, since it depends on the motion parameter (see Suppl.
+Mat.
+of [29]).
+Although our denoising efficacy is confirmed from multiple perspectives, we believe that advancing denoising metrics for real-world data (where it gets interesting due to the absence of GT labels) is a central direction for future work in event-based vision.
 6 Conclusion
-Motion is an intrinsic part of event camera data, since scene edges cannot be sensed without it.
-We propose the first joint estimation framework for denoising and generic motion estimation, which connects two topics that have traditionally been disconnected in event-based vision.
-The experiments show that the proposed method achieves state-of-the-art results on the E-MLB benchmark and competitive results on the DND21 benchmark for denoising, while showing its efficacy on motion estimation and intensity reconstruction tasks.
-Moreover, the proposed method is flexible, as it allows replacing the 1-step motion estimation of CMax with any other motion estimator, such as DNNs.
-We believe that the proposed approach contributes not only to strengthening the theory of event-data denoising but also to expanding practical denoising use-cases via the code published.
+In summary, this work proposes that: ($i$) Joint estimation of noise and motion is a novel problem setting compared to conventional (isolated) denoising, and is sensitive given the event data’s inherent motion dependency.
+($ii$) The proposed joint estimation framework can accommodate various motion estimators and has shown its efficacy in improving both denoising and motion estimation accuracies.
+($iii$) It is important to evaluate denoising results on both, denoising benchmarks and downstream tasks (e.g., motion estimation, intensity reconstruction), since current denoising benchmarks have shortcomings in terms of metrics, data, and baselines.
+In particular, our method achieves notable denoising results (state-of-the-art on the E-MLB benchmark and competitive on DND21) while demonstrating effectiveness across motion estimation and intensity reconstruction tasks.
+Our approach advances both event-denoising theory and practical denoising use-cases.
 Acknowledgments
@@
 Jiang for useful discussions.
-This research was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2002/1 “Science of Intelligence” – project number 390523135.
+Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2002/1 “Science of Intelligence” – project number 390523135.
 7 Supplementary
 7.1 Video
-We encourage readers to inspect the attached video, which summarizes the method and the results.
-$\text{iter}=0$
+We encourage readers to inspect the attached video, which summarizes the method and results.
+[…]
 Refer to caption Refer to caption Refer to caption
-$\text{iter}=20$
+[…]
 Refer to caption Refer to caption Refer to caption
-$\text{iter}=100$
+[…]
 Refer to caption Refer to caption Refer to caption
@@
 Figure 11: Evolutions of signal, noise, and motion (IWE) during optimization.
-The edge structure (e.g., green boxes in (b) Noise) converges to move to signal events, while the CMax converges to the sharp IWE (i.e., expected motion parameters).
+The edge structure (e.g., green boxes in (b) Noise) converges to move to signal events, while CMax converges to the sharp IWE (i.e., expected motion parameters).
 (a) Signal (first iter)
@@
 Figure 12: Results on HD real-world data and intermediate denoising values (TPR, FPR and ESR metrics) during optimization.
-1Hz 3Hz 5Hz 7Hz 10Hz hotel driving hotel driving hotel driving hotel driving hotel driving Model-based KNoise [30] $0.6773$ $0.6296$ $0.6521$ $0.6230$ $0.6703$ $0.6235$ $0.6583$ $0.6164$ $0.6413$ $0.6142$ DWF [24] $0.9268$ $0.7409$ $0.8930$ $0.7099$ $0.8620$ $0.6901$ $0.8338$ $0.6747$ $0.7958$ $0.6563$ BAF [9] $0.9535$ $0.8479$ $0.9197$ $0.8155$ $0.8916$ $0.7930$ $0.8662$ $0.7732$ $0.8366$ $0.7479$ Ynoise [14] $0.9690$ $0.9409$ $0.9517$ $0.9240$ $0.9234$ $0.9093$ $0.9177$ $0.8972$ $0.8987$ $0.8800$ TS [33] $0.9716$ $0.9307$ $0.9721$ $0.9260$ $0.9606$ $0.9270$ $0.9654$ $0.9241$ $0.9620$ $0.9202$ Ours 1.013 802 194 317 176 21.01380219431717621.013\,802\,194\,317\,176\,21.013 802 194 317 176 2 0.882 119 693 133 694 80.88211969313369480.882\,119\,693\,133\,694\,80.882 119 693 133 694 8 0.967 937 633 567 141 90.96793763356714190.967\,937\,633\,567\,141\,90.967 937 633 567 141 9 0.850 838 015 623 797 50.85083801562379750.850\,838\,015\,623\,797\,50.850 838 015 623 797 5 0.962 515 815 619 444 40.96251581561944440.962\,515\,815\,619\,444\,40.962 515 815 619 444 4 0.855 398 484 638 116 20.85539848463811620.855\,398\,484\,638\,116\,20.855 398 484 638 116 2 0.951 406 776 235 535 30.95140677623553530.951\,406\,776\,235\,535\,30.951 406 776 235 535 3 0.846 774 906 545 3740.8467749065453740.846\,774\,906\,545\,3740.846 774 906 545 374 0.960 564 724 999 644 20.96056472499964420.960\,564\,724\,999\,644\,20.960 564 724 999 644 2 0.836 273 555 102 2360.8362735551022360.836\,273\,555\,102\,2360.836 273 555 102 236 Learning EDnCNN [4] $0.9573$ $0.8873$ $0.9371$ $0.8771$ $0.9365$ $0.8748$ $0.9254$ $0.8654$ $0.9006$ $0.874$ MLPF [24] $0.9704$ $0.8887$ $0.9718$ $0.8873$ $0.9704$ $0.8845$ $0.9691$ $0.8817$ $0.9634$ $0.8761$ EDformer [28] $0.9928$ $0.9541$ $0.9891$ $0.9472$ $0.9845$ $0.9424$ $0.9792$ $0.9343$ $0.9699$ $0.9264$
+1Hz 3Hz 5Hz 7Hz 10Hz hotel driving hotel driving hotel driving hotel driving hotel driving Model-based BAF [9] 0.95350.95350.9535 0.84790.84790.8479 0.91970.91970.9197 0.81550.81550.8155 0.89160.89160.8916 0.79300.79300.7930 0.86620.86620.8662 0.77320.77320.7732 0.83660.83660.8366 0.74790.74790.7479 TS [34] 0.97160.97160.9716 0.93070.93070.9307 0.97210.97210.9721 0.92600.92600.9260 0.96060.96060.9606 0.92700.92700.9270 0.96540.96540.9654 0.92410.92410.9241 0.96200.96200.9620 0.92020.92020.9202 KNoise [31] 0.67730.67730.6773 0.62960.62960.6296 0.65210.65210.6521 0.62300.62300.6230 0.67030.67030.6703 0.62350.62350.6235 0.65830.65830.6583 0.61640.61640.6164 0.64130.64130.6413 0.61420.61420.6142 Ynoise [14] 0.96900.96900.9690 0.94090.94090.9409 0.95170.95170.9517 0.92400.92400.9240 0.92340.92340.9234 0.90930.90930.9093 0.91770.91770.9177 0.89720.89720.8972 0.89870.89870.8987 0.88000.88000.8800 DWF [24] 0.92680.92680.9268 0.74090.74090.7409 0.89300.89300.8930 0.70990.70990.7099 0.86200.86200.8620 0.69010.69010.6901 0.83380.83380.8338 0.67470.67470.6747 0.79580.79580.7958 0.65630.65630.6563 Ours 1.013 802 194 317 176 21.013\,802\,194\,317\,176\,21.013 802 194 317 176 2 0.882 119 693 133 694 80.882\,119\,693\,133\,694\,80.882 119 693 133 694 8 0.967 937 633 567 141 90.967\,937\,633\,567\,141\,90.967 937 633 567 141 9 0.850 838 015 623 797 50.850\,838\,015\,623\,797\,50.850 838 015 623 797 5 0.962 515 815 619 444 40.962\,515\,815\,619\,444\,40.962 515 815 619 444 4 0.855 398 484 638 116 20.855\,398\,484\,638\,116\,20.855 398 484 638 116 2 0.951 406 776 235 535 30.951\,406\,776\,235\,535\,30.951 406 776 235 535 3 0.846 774 906 545 3740.846\,774\,906\,545\,3740.846 774 906 545 374 0.960 564 724 999 644 20.960\,564\,724\,999\,644\,20.960 564 724 999 644 2 0.836 273 555 102 2360.836\,273\,555\,102\,2360.836 273 555 102 236 Learning EDnCNN [4] 0.95730.95730.9573 0.88730.88730.8873 0.93710.93710.9371 0.87710.87710.8771 0.93650.93650.9365 0.87480.87480.8748 0.92540.92540.9254 0.86540.86540.8654 0.90060.90060.9006 0.8740.8740.874 MLPF [24] 0.97040.97040.9704 0.88870.88870.8887 0.97180.97180.9718 0.88730.88730.8873 0.97040.97040.9704 0.88450.88450.8845 0.96910.96910.9691 0.88170.88170.8817 0.96340.96340.9634 0.87610.87610.8761 EDformer [29] 0.99280.99280.9928 0.95410.95410.9541 0.98910.98910.9891 0.94720.94720.9472 0.98450.98450.9845 0.94240.94240.9424 0.97920.97920.9792 0.93430.93430.9343 0.96990.96990.9699 0.92640.92640.9264
 […]
-dynamic_rot boxes_rot RMS $\downarrow$ FWL $\uparrow$ RMS $\downarrow$ FWL $\uparrow$ $5$ Hz Noise CMax [17] 20.000 709 6820.0007096820.000\,709\,6820.000 709 68 1.258 796 0971.2587960971.258\,796\,0971.258 796 097 124.641 377124.641377124.641\,377124.641 377 1.128 529 5931.1285295931.128\,529\,5931.128 529 593 – w/ Init.
-8.274 876 2458.2748762458.274\,876\,2458.274 876 245 1.273 697 0451.2736970451.273\,697\,0451.273 697 045 20.659 240 3320.6592403320.659\,240\,3320.659 240 33 1.214 159 4591.2141594591.214\,159\,4591.214 159 459 Downsampling $90$% 8.807 703 958.807703958.807\,703\,958.807 703 95 1.272 907 8221.2729078221.272\,907\,8221.272 907 822 124.868 639 7124.8686397124.868\,639\,7124.868 639 7 1.107 758 3951.1077583951.107\,758\,3951.107 758 395 – w/ Init.
-8.226 219 4448.2262194448.226\,219\,4448.226 219 444 1.273 725 3071.2737253071.273\,725\,3071.273 725 307 20.618 615 8620.6186158620.618\,615\,8620.618 615 86 1.214 188 4551.2141884551.214\,188\,4551.214 188 455 Downsampling $80$% 9.964 567 2999.9645672999.964\,567\,2999.964 567 299 1.270 984 5391.2709845391.270\,984\,5391.270 984 539 161.810 797 8161.8107978161.810\,797\,8161.810 797 8 1.060 671 $1$.0606711111.060\,671\,$1$.060 671 111 – w/ Init.
-8.243 565 3688.2435653688.243\,565\,3688.243 565 368 1.273 741 8421.2737418421.273\,741\,8421.273 741 842 20.797 769 5620.7977695620.797\,769\,5620.797 769 56 1.214 169 4341.2141694341.214\,169\,4341.214 169 434 Downsampling $70$% 14.398 878 9114.3988789114.398\,878\,9114.398 878 91 1.262 100 3441.2621003441.262\,100\,3441.262 100 344 184.511 383 4184.5113834184.511\,383\,4184.511 383 4 1.030 485 5451.0304855451.030\,485\,5451.030 485 545 – w/ Init.
-8.231 444 7738.2314447738.231\,444\,7738.231 444 773 1.273 783 9431.2737839431.273\,783\,9431.273 783 943 23.678 913 0923.6789130923.678\,913\,0923.678 913 09 1.210 643 261.210643261.210\,643\,261.210 643 26 Ours $90$% 8.521 574 0738.5215740738.521\,574\,0738.521 574 073 1.273 446 2961.2734462961.273\,446\,2961.273 446 296 97.584 574 8497.5845748497.584\,574\,8497.584 574 84 1.140 880 661.140880661.140\,880\,661.140 880 66 – w/ Init.
-8.188 607 2268.1886072268.188\,607\,2268.188 607 226 1.273 725 2991.2737252991.273\,725\,2991.273 725 299 20.604 334 6220.6043346220.604\,334\,6220.604 334 62 1.214 179 0181.2141790181.214\,179\,0181.214 179 018 Ours $80$% 8.511 006 6718.5110066718.511\,006\,6718.511 006 671 1.273 293 7311.2732937311.273\,293\,7311.273 293 731 131.355 614 8131.3556148131.355\,614\,8131.355 614 8 1.102 904 5081.1029045081.102\,904\,5081.102 904 508 – w/ Init.
-8.170 039 698.170039698.170\,039\,698.170 039 69 1.273 790 7891.2737907891.273\,790\,7891.273 790 789 20.861 568 0320.8615680320.861\,568\,0320.861 568 03 1.214 094 851.214094851.214\,094\,851.214 094 85 Ours $70$% 9.180 407 8239.1804078239.180\,407\,8239.180 407 823 1.272 074 7881.2720747881.272\,074\,7881.272 074 788 163.676 240 5163.6762405163.676\,240\,5163.676 240 5 1.061 940 9851.0619409851.061\,940\,9851.061 940 985 – w/ Init.
-8.086 262 248.086262248.086\,262\,248.086 262 24 1.273 578 8461.2735788461.273\,578\,8461.273 578 846 21.151 118 9621.1511189621.151\,118\,9621.151 118 96 1.213 969 0641.2139690641.213\,969\,0641.213 969 064 BAF 19.674 893 0519.6748930519.674\,893\,0519.674 893 05 1.259 772 491.259772491.259\,772\,491.259 772 49 125.028 145 1125.0281451125.028\,145\,1125.028 145 1 1.127 308 9021.1273089021.127\,308\,9021.127 308 902 – w/ Init.
-8.253 078 2028.2530782028.253\,078\,2028.253 078 202 1.273 636 991.273636991.273\,636\,991.273 636 99 19.550 436 619.550436619.550\,436\,619.550 436 6 1.213 505 481.213505481.213\,505\,481.213 505 48 $1$ Hz CMax [17] 19.395 190 2319.3951902319.395\,190\,2319.395 190 23 1.275 866 1481.2758661481.275\,866\,1481.275 866 148 117.439 798 6117.4397986117.439\,798\,6117.439 798 6 1.143 852 7451.1438527451.143\,852\,7451.143 852 745 – w/ Init.
-8.254 263 2638.2542632638.254\,263\,2638.254 263 263 1.289 966 6551.2899666551.289\,966\,6551.289 966 655 20.627 507 7120.6275077120.627\,507\,7120.627 507 71 1.223 032 6881.2230326881.223\,032\,6881.223 032 688 Downsampling $90$% 8.676 346 0658.6763460658.676\,346\,0658.676 346 065 1.289 459 4391.2894594391.289\,459\,4391.289 459 439 110.568 366 3110.5683663110.568\,366\,3110.568 366 3 1.130 495 7731.1304957731.130\,495\,7731.130 495 773 – w/ Init.
-8.184 187 8618.1841878618.184\,187\,8618.184 187 861 1.289 611 9121.2896119121.289\,611\,9121.289 611 912 20.620 358 2120.6203582120.620\,358\,2120.620 358 21 1.223 050 3321.2230503321.223\,050\,3321.223 050 332 Ours $90$% 8.505 807 8288.5058078288.505\,807\,8288.505 807 828 1.289 787 0891.2897870891.289\,787\,0891.289 787 089 87.774 541 8687.7745418687.774\,541\,8687.774 541 86 1.158 678 1341.1586781341.158\,678\,1341.158 678 134 – w/ Init.
-8.177 036 3548.1770363548.177\,036\,3548.177 036 354 1.290 007 1851.2900071851.290\,007\,1851.290 007 185 20.569 191 6120.5691916120.569\,191\,6120.569 191 61 1.223 052 861.223052861.223\,052\,861.223 052 86 BAF 19.568 905 3719.5689053719.568\,905\,3719.568 905 37 1.275 738 1221.2757381221.275\,738\,1221.275 738 122 117.554 331 4117.5543314117.554\,331\,4117.554 331 4 1.143 238 0781.1432380781.143\,238\,0781.143 238 078 – w/ Init.
-8.189 429 3938.1894293938.189\,429\,3938.189 429 393 1.289 532 0351.2895320351.289\,532\,0351.289 532 035 20.713 437 520.713437520.713\,437\,520.713 437 5 1.223 006 3741.2230063741.223\,006\,3741.223 006 374
+dynamic_rot boxes_rot RMS $\downarrow$ FWL $\uparrow$ RMS $\downarrow$ FWL $\uparrow$ 555 Hz Noise CMax [17] 20.000 709 6820.000\,709\,6820.000 709 68 1.258 796 0971.258\,796\,0971.258 796 097 124.641 377124.641\,377124.641 377 1.128 529 5931.128\,529\,5931.128 529 593 – w/ Init.
+8.274 876 2458.274\,876\,2458.274 876 245 1.273 697 0451.273\,697\,0451.273 697 045 20.659 240 3320.659\,240\,3320.659 240 33 1.214 159 4591.214\,159\,4591.214 159 459 Downsampling 909090% 8.807 703 958.807\,703\,958.807 703 95 1.272 907 8221.272\,907\,8221.272 907 822 124.868 639 7124.868\,639\,7124.868 639 7 1.107 758 3951.107\,758\,3951.107 758 395 – w/ Init.
+8.226 219 4448.226\,219\,4448.226 219 444 1.273 725 3071.273\,725\,3071.273 725 307 20.618 615 8620.618\,615\,8620.618 615 86 1.214 188 4551.214\,188\,4551.214 188 455 Downsampling 808080% 9.964 567 2999.964\,567\,2999.964 567 299 1.270 984 5391.270\,984\,5391.270 984 539 161.810 797 8161.810\,797\,8161.810 797 8 1.060 671 $1$.060\,671\,$1$.060 671 111 – w/ Init.
+8.243 565 3688.243\,565\,3688.243 565 368 1.273 741 8421.273\,741\,8421.273 741 842 20.797 769 5620.797\,769\,5620.797 769 56 1.214 169 4341.214\,169\,4341.214 169 434 Downsampling 707070% 14.398 878 9114.398\,878\,9114.398 878 91 1.262 100 3441.262\,100\,3441.262 100 344 184.511 383 4184.511\,383\,4184.511 383 4 1.030 485 5451.030\,485\,5451.030 485 545 – w/ Init.
+8.231 444 7738.231\,444\,7738.231 444 773 1.273 783 9431.273\,783\,9431.273 783 943 23.678 913 0923.678\,913\,0923.678 913 09 1.210 643 261.210\,643\,261.210 643 26 Ours 909090% 8.521 574 0738.521\,574\,0738.521 574 073 1.273 446 2961.273\,446\,2961.273 446 296 97.584 574 8497.584\,574\,8497.584 574 84 1.140 880 661.140\,880\,661.140 880 66 – w/ Init.
+8.188 607 2268.188\,607\,2268.188 607 226 1.273 725 2991.273\,725\,2991.273 725 299 20.604 334 6220.604\,334\,6220.604 334 62 1.214 179 0181.214\,179\,0181.214 179 018 Ours 808080% 8.511 006 6718.511\,006\,6718.511 006 671 1.273 293 7311.273\,293\,7311.273 293 731 131.355 614 8131.355\,614\,8131.355 614 8 1.102 904 5081.102\,904\,5081.102 904 508 – w/ Init.
+8.170 039 698.170\,039\,698.170 039 69 1.273 790 7891.273\,790\,7891.273 790 789 20.861 568 0320.861\,568\,0320.861 568 03 1.214 094 851.214\,094\,851.214 094 85 Ours 707070% 9.180 407 8239.180\,407\,8239.180 407 823 1.272 074 7881.272\,074\,7881.272 074 788 163.676 240 5163.676\,240\,5163.676 240 5 1.061 940 9851.061\,940\,9851.061 940 985 – w/ Init.
+8.086 262 248.086\,262\,248.086 262 24 1.273 578 8461.273\,578\,8461.273 578 846 21.151 118 9621.151\,118\,9621.151 118 96 1.213 969 0641.213\,969\,0641.213 969 064 BAF 19.674 893 0519.674\,893\,0519.674 893 05 1.259 772 491.259\,772\,491.259 772 49 125.028 145 1125.028\,145\,1125.028 145 1 1.127 308 9021.127\,308\,9021.127 308 902 – w/ Init.
+8.253 078 2028.253\,078\,2028.253 078 202 1.273 636 991.273\,636\,991.273 636 99 19.550 436 619.550\,436\,619.550 436 6 1.213 505 481.213\,505\,481.213 505 48 111 Hz CMax [17] 19.395 190 2319.395\,190\,2319.395 190 23 1.275 866 1481.275\,866\,1481.275 866 148 117.439 798 6117.439\,798\,6117.439 798 6 1.143 852 7451.143\,852\,7451.143 852 745 – w/ Init.
+8.254 263 2638.254\,263\,2638.254 263 263 1.289 966 6551.289\,966\,6551.289 966 655 20.627 507 7120.627\,507\,7120.627 507 71 1.223 032 6881.223\,032\,6881.223 032 688 Downsampling 909090% 8.676 346 0658.676\,346\,0658.676 346 065 1.289 459 4391.289\,459\,4391.289 459 439 110.568 366 3110.568\,366\,3110.568 366 3 1.130 495 7731.130\,495\,7731.130 495 773 – w/ Init.
+8.184 187 8618.184\,187\,8618.184 187 861 1.289 611 9121.289\,611\,9121.289 611 912 20.620 358 2120.620\,358\,2120.620 358 21 1.223 050 3321.223\,050\,3321.223 050 332 Ours 909090% 8.505 807 8288.505\,807\,8288.505 807 828 1.289 787 0891.289\,787\,0891.289 787 089 87.774 541 8687.774\,541\,8687.774 541 86 1.158 678 1341.158\,678\,1341.158 678 134 – w/ Init.
+8.177 036 3548.177\,036\,3548.177 036 354 1.290 007 1851.290\,007\,1851.290 007 185 20.569 191 6120.569\,191\,6120.569 191 61 1.223 052 861.223\,052\,861.223 052 86 BAF 19.568 905 3719.568\,905\,3719.568 905 37 1.275 738 1221.275\,738\,1221.275 738 122 117.554 331 4117.554\,331\,4117.554 331 4 1.143 238 0781.143\,238\,0781.143 238 078 – w/ Init.
+8.189 429 3938.189\,429\,3938.189 429 393 1.289 532 0351.289\,532\,0351.289 532 035 20.713 437 520.713\,437\,520.713 437 5 1.223 006 3741.223\,006\,3741.223 006 374
 Table 6: Angular velocity estimation on ECD dataset [40].
+7.2 Experiment details
+Hyper-parameters.
+For the denoising experiments, we test various values $\tau=\{0.9,\ldots,0.1\}$ to calculate the ROC on DND21 and $\tau=\{0.9,\ldots,0.7\}$ for the RMS on ECD.
+For E-MLB benchmarking, we fix the number of signal events to follow the prior work [10].
+To analyze the proposed pipeline in the CMax framework, we use model-based rotational motion estimation [15] and tile-based optical flow estimation [48] approaches.
+We use the magnitude of the IWE gradient [18] as the CMax objective function.
 7.3 Denoising Convergence During Optimization
@@
 5).
-At the first iteration (i.e., initialization), signal and noise events are classified randomly.
-As the optimization proceeds, signal events evolve to keep edge structures in the scene, while noise events evolve to drop such edge structures (see the second and third rows).
+In the first iteration (i.e., initialization), signal and noise events are randomly classified.
+As optimization proceeds, signal events evolve towards keeping edge structures in the scene, while noise events evolve towards dropping such edge structures (see second and third rows).
 Also, the IWEs converge to sharp edges with correct motion parameters.
-This example further confirms the efficacy of the joint estimation.
-The optimization process on an HD-resolution ($1280\times 720$ px) real-world dataset, TUM-VIE [32], is shown in Fig.
+This example confirms the efficacy of joint estimation.
+The optimization process on an HD-resolution (1280×7201280$\times$ 7201280 × 720 px) real-world dataset, TUM-VIE [33], is shown in Fig.
 12.
@@
 The original CMax degrades due to noise, as reported in previous work (e.g., [3]).
-The S/N target ratio $\tau$ affects the accuracy: when it is close to the actual value of noise injection, the result of the proposed method becomes better.
-The amount of artificial noise injected is around $15$ for $5$ Hz and $3$ for $1$ Hz conditions.
-Although we cannot know the “true” noise level due to the original noise in the ECD sequences, our method constantly produces better accuracy and FWL than other baselines.
+The signal-to-noise (S/N) target ratio $\tau$ affects accuracy: the closer it is to the actual value of noise injection, the better the results of the proposed method.
+The amount of artificial noise injected is around 151515 % for 555 Hz and 333 % for 111 Hz conditions.
+Although the “true” noise level is unknown due to the original noise in the ECD sequences, our method constantly produces better accuracy and FWL values than the baselines.
 Please refer to Sec.
-4.3 for more discussions about dependency on initialization and comparison with other baselines.
-The AUCs for the conditions that we test ($\tau=\{0.9,\ldots,0.7\}$) are $0.70$ (“Ours”) and $0.67$ (“Downsampling”).
+4.3 for more discussions about the dependency on initialization and comparison with other baselines.
+The AUCs for the conditions that we test ($\tau=\{0.9,\ldots,0.7\}$) are 0.700.700.70 (“Ours”) and 0.670.670.67 (“Downsampling”).
 7.6 Quantitative Evaluation of Intensity Reconstruction
@@
 2 and 7: the frames become underexposed or blurry due to their limited dynamic range, when event data suffer from more BA noise (i.e., in dark scenes).
-Nonetheless, we report non-reference image quality indices for different S/N ratios ($\tau$).
+Nonetheless, we report non-reference image quality indices for different S/N ratios $\tau$.
 Figure 13 reports the scores of Blind/Referenceless Image Spatial Quality Evaluator (BRISQUE) [38] and Naturalness Image Quality Evaluator (NIQE) [39], using Bicycle-ND64-2 sequence (same as Fig.
@@
 Although the results potentially suggest that it could estimate the “true” noise ratio in the data using the non-reference indices, which is useful for image reconstruction applications, we leave further evaluation and discussion as future work.
-Figure 13: Results of the non-reference image quality indices for image reconstruction. See Fig. 8 for the details.
-7 Limitations
-Our method expects as parameter a target noise level, but the intrinsic (“true”) noise level depends on the camera, scene, etc.
-and is therefore unknown.
-More research in this direction is needed for automatic parameter tuning.
-Also, notice that there might be a potential limitation of the denoising metric ESR, since it depends on the motion parameter (see Suppl.
-Mat.
-of [28]).
-Although our denoising efficacy is confirmed from multiple perspectives, we believe that advancing denoising metrics for real-world data (where it gets interesting, as there are no GT labels) is a central direction for future work in event-based vision.
+Figure 13: Results of the non-reference image quality indices for image reconstruction. See Fig. 8 for details.